### v1

# A. Abed Abud CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland B. Abi

A. Abed Abud
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
B. Abi
University of Oxford, Oxford, OX1 3RH, United Kingdom
R. Acciarri
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
M. A. Acero
Universidad del Atlántico, Barranquilla, Atlántico, Colombia
M. R. Adames
Universidade Tecnológica Federal do Paraná, Curitiba, Brazil
G. Adamov
Georgian Technical University, Tbilisi, Georgia
M. Adamowski
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
D. Adams
Brookhaven National Laboratory, Upton, NY 11973, USA
M. Adinolfi
University of Bristol, Bristol BS8 1TL, United Kingdom
C. Adriano
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
A. Aduszkiewicz
University of Houston, Houston, TX 77204, USA
J. Aguilar
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
B. Aimard
Laboratoire d’Annecy de Physique des Particules, Université Savoie Mont Blanc, CNRS, LAPP-IN2P3, 74000 Annecy, France
F. Akbar
University of Rochester, Rochester, NY 14627, USA
K. Allison
University of Colorado Boulder, Boulder, CO 80309, USA
S. Alonso Monsalve
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
M. Alrashed
Kansas State University, Manhattan, KS 66506, USA
A. Alton
Augustana University, Sioux Falls, SD 57197, USA
R. Alvarez
CIEMAT, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas, E-28040 Madrid, Spain
T. Alves
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
H. Amar
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
P. Amedo
Instituto Galego de Física de Altas Enerxías, University of Santiago de Compostela, Santiago de Compostela, 15782, Spain
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
J. Anderson
Argonne National Laboratory, Argonne, IL 60439, USA
D. A. Andrade
Illinois Institute of Technology, Chicago, IL 60616, USA
C. Andreopoulos
University of Liverpool, L69 7ZE, Liverpool, United Kingdom
M. Andreotti
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
University of Ferrara, Ferrara, Italy
M. P. Andrews
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
F. Andrianala
University of Antananarivo, Antananarivo 101, Madagascar
S. Andringa
Laboratório de Instrumentação e Física Experimental de Partículas, 1649-003 Lisboa and 3004-516 Coimbra, Portugal
[…]
A. Ankowski
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
M. Antoniassi
Universidade Tecnológica Federal do Paraná, Curitiba, Brazil
M. Antonova
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
[…]
A. Aranda-Fernandez
Universidad de Colima, Colima, Mexico
L. Arellano
University of Manchester, Manchester M13 9PL, United Kingdom
E. Arrieta Diaz
Universidad del Magdalena, Santa Marta - Colombia
M. A. Arroyave
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
J. Asaadi
University of Texas at Arlington, Arlington, TX 76019, USA
A. Ashkenazi
Tel Aviv University, Tel Aviv-Yafo, Israel
D. Asner
Brookhaven National Laboratory, Upton, NY 11973, USA
L. Asquith
University of Sussex, Brighton, BN1 9RH, United Kingdom
E. Atkin
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
D. Auguste
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
A. Aurisano
University of Cincinnati, Cincinnati, OH 45221, USA
V. Aushev
Taras Shevchenko National University of Kyiv, 01601 Kyiv, Ukraine
D. Autiero
Institut de Physique des 2 Infinis de Lyon, 69622 Villeurbanne, France
F. Azfar
University of Oxford, Oxford, OX1 3RH, United Kingdom
A. Back
Indiana University, Bloomington, IN 47405, USA
H. Back
Pacific Northwest National Laboratory, Richland, WA 99352, USA
J. J. Back
University of Warwick, Coventry CV4 7AL, United Kingdom
I. Bagaturia
Georgian Technical University, Tbilisi, Georgia
L. Bagby
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
[…]
S. Balasubramanian
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
P. Baldi
University of California Irvine, Irvine, CA 92697, USA
W. Baldini
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
J. Baldonedo
University of Vigo, E- 36310 Vigo Spain
B. Baller
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
B. Bambah
University of Hyderabad, Gachibowli, Hyderabad - 500 046, India
R. Banerjee
York University, Toronto M3J 1P3, Canada
F. Barao
Laboratório de Instrumentação e Física Experimental de Partículas, 1649-003 Lisboa and 3004-516 Coimbra, Portugal
Instituto Superior Técnico - IST, Universidade de Lisboa, Portugal
G. Barenboim
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
P. Barham Alzás
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
G. J. Barker
University of Warwick, Coventry CV4 7AL, United Kingdom
W. Barkhouse
University of North Dakota, Grand Forks, ND 58202-8357, USA
G. Barr
University of Oxford, Oxford, OX1 3RH, United Kingdom
J. Barranco Monarca
Universidad de Guanajuato, Guanajuato, C.P. 37000, Mexico
A. Barros
Universidade Tecnológica Federal do Paraná, Curitiba, Brazil
N. Barros
Laboratório de Instrumentação e Física Experimental de Partículas, 1649-003 Lisboa and 3004-516 Coimbra, Portugal
Faculdade de Ciências da Universidade de Lisboa - FCUL, 1749-016 Lisboa, Portugal
D. Barrow
University of Oxford, Oxford, OX1 3RH, United Kingdom
J. L. Barrow
Massachusetts Institute of Technology, Cambridge, MA 02139, USA
A. Basharina-Freshville
University College London, London, WC1E 6BT, United Kingdom
A. Bashyal
Argonne National Laboratory, Argonne, IL 60439, USA
V. Basque
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
C. Batchelor
University of Edinburgh, Edinburgh EH8 9YL, United Kingdom
L. Bathe-Peters
University of Oxford, Oxford, OX1 3RH, United Kingdom
J.B.R. Battat
Wellesley College, Wellesley, MA 02481, USA
F. Battisti
University of Oxford, Oxford, OX1 3RH, United Kingdom
F. Bay
Antalya Bilim University, 07190 Döşemealtı/Antalya, Turkey
M. C. Q. Bazetto
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
J. L. L. Bazo Alba
Pontificia Universidad Católica del Perú, Lima, Perú
J. F. Beacom
Ohio State University, Columbus, OH 43210, USA
E. Bechetoille
Institut de Physique des 2 Infinis de Lyon, 69622 Villeurbanne, France
B. Behera
University of Florida, Gainesville, FL 32611-8440, USA
E. Belchior
Louisiana State University, Baton Rouge, LA 70803, USA
G. Bell
Daresbury Laboratory, Cheshire WA4 4AD, United Kingdom
L. Bellantoni
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
G. Bellettini
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
Università di Pisa, I-56127 Pisa, Italy
V. Bellini
Istituto Nazionale di Fisica Nucleare Sezione di Catania, I-95123 Catania, Italy
Università di Catania, 2 - 95131 Catania, Italy
O. Beltramello
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
N. Benekos
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
C. Benitez Montiel
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
Universidad Nacional de Asunción, San Lorenzo, Paraguay
D. Benjamin
Brookhaven National Laboratory, Upton, NY 11973, USA
F. Bento Neves
Laboratório de Instrumentação e Física Experimental de Partículas, 1649-003 Lisboa and 3004-516 Coimbra, Portugal
J. Berger
Colorado State University, Fort Collins, CO 80523, USA
S. Berkman
Michigan State University, East Lansing, MI 48824, USA
J. Bernal
Universidad Nacional de Asunción, San Lorenzo, Paraguay
P. Bernardini
Istituto Nazionale di Fisica Nucleare Sezione di Lecce, 73100 - Lecce, Italy
Università del Salento, 73100 Lecce, Italy
A. Bersani
Istituto Nazionale di Fisica Nucleare Sezione di Genova, 16146 Genova GE, Italy
S. Bertolucci
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
M. Betancourt
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Betancur Rodríguez
Universidad EIA, Envigado, Antioquia, Colombia
A. Bevan
Queen Mary University of London, London E1 4NS, United Kingdom
Y. Bezawada
University of California Davis, Davis, CA 95616, USA
A. T. Bezerra
Universidade Federal de Alfenas, Poços de Caldas - MG, 37715-400, Brazil
T. J. Bezerra
University of Sussex, Brighton, BN1 9RH, United Kingdom
A. Bhat
University of Chicago, Chicago, IL 60637, USA
V. Bhatnagar
Panjab University, Chandigarh, 160014, India
J. Bhatt
University College London, London, WC1E 6BT, United Kingdom
M. Bhattacharjee
Indian Institute of Technology Guwahati, Guwahati, 781 039, India
M. Bhattacharya
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
S. Bhuller
University of Bristol, Bristol BS8 1TL, United Kingdom
B. Bhuyan
Indian Institute of Technology Guwahati, Guwahati, 781 039, India
S. Biagi
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali del Sud, 95123 Catania, Italy
J. Bian
University of California Irvine, Irvine, CA 92697, USA
K. Biery
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
B. Bilki
Beykent University, Istanbul, Turkey
University of Iowa, Iowa City, IA 52242, USA
M. Bishai
Brookhaven National Laboratory, Upton, NY 11973, USA
A. Bitadze
University of Manchester, Manchester M13 9PL, United Kingdom
A. Blake
Lancaster University, Lancaster LA1 4YB, United Kingdom
F. D. Blaszczyk
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
G. C. Blazey
Northern Illinois University, DeKalb, IL 60115, USA
E. Blucher
University of Chicago, Chicago, IL 60637, USA
J. Bogenschuetz
University of Texas at Arlington, Arlington, TX 76019, USA
J. Boissevain
Los Alamos National Laboratory, Los Alamos, NM 87545, USA
S. Bolognesi
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
T. Bolton
Kansas State University, Manhattan, KS 66506, USA
L. Bomben
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
University of Insubria, Via Ravasi, 2, 21100 Varese VA, Italy
M. Bonesini
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
Università di Milano Bicocca , 20126 Milano, Italy
C. Bonilla-Diaz
Universidad Católica del Norte, Antofagasta, Chile
F. Bonini
Brookhaven National Laboratory, Upton, NY 11973, USA
A. Booth
Queen Mary University of London, London E1 4NS, United Kingdom
F. Boran
Indiana University, Bloomington, IN 47405, USA
S. Bordoni
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
R. Borges Merlo
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
A. Borkum
University of Sussex, Brighton, BN1 9RH, United Kingdom
N. Bostan
University of Iowa, Iowa City, IA 52242, USA
J. Bracinik
University of Birmingham, Birmingham B15 2TT, United Kingdom
D. Braga
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
B. Brahma
Indian Institute of Technology Hyderabad, Hyderabad, 502285, India
D. Brailsford
Lancaster University, Lancaster LA1 4YB, United Kingdom
F. Bramati
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
A. Branca
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
A. Brandt
University of Texas at Arlington, Arlington, TX 76019, USA
J. Bremer
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
C. Brew
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
S. J. Brice
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
V. Brio
Istituto Nazionale di Fisica Nucleare Sezione di Catania, I-95123 Catania, Italy
C. Brizzolari
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
Università di Milano Bicocca , 20126 Milano, Italy
C. Bromberg
Michigan State University, East Lansing, MI 48824, USA
J. Brooke
University of Bristol, Bristol BS8 1TL, United Kingdom
A. Bross
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
G. Brunetti
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
Università di Milano Bicocca , 20126 Milano, Italy
M. Brunetti
University of Warwick, Coventry CV4 7AL, United Kingdom
N. Buchanan
Colorado State University, Fort Collins, CO 80523, USA
H. Budd
University of Rochester, Rochester, NY 14627, USA
J. Buergi
University of Bern, CH-3012 Bern, Switzerland
D. Burgardt
Wichita State University, Wichita, KS 67260, USA
S. Butchart
University of Sussex, Brighton, BN1 9RH, United Kingdom
G. Caceres V
University of California Davis, Davis, CA 95616, USA
I. Cagnoli
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
T. Cai
York University, Toronto M3J 1P3, Canada
R. Calabrese
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
University of Ferrara, Ferrara, Italy
J. Calcutt
Oregon State University, Corvallis, OR 97331, USA
M. Calin
University of Bucharest, Bucharest, Romania
L. Calivers
University of Bern, CH-3012 Bern, Switzerland
E. Calvo
CIEMAT, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas, E-28040 Madrid, Spain
A. Caminata
Istituto Nazionale di Fisica Nucleare Sezione di Genova, 16146 Genova GE, Italy
A. F. Camino
University of Pittsburgh, Pittsburgh, PA 15260, USA
W. Campanelli
Laboratório de Instrumentação e Física Experimental de Partículas, 1649-003 Lisboa and 3004-516 Coimbra, Portugal
A. Campani
Istituto Nazionale di Fisica Nucleare Sezione di Genova, 16146 Genova GE, Italy
Università degli Studi di Genova, Genova, Italy
A. Campos Benitez
Virginia Tech, Blacksburg, VA 24060, USA
N. Canci
Istituto Nazionale di Fisica Nucleare Sezione di Napoli, I-80126 Napoli, Italy
J. Capó
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
I. Caracas
Johannes Gutenberg-Universität Mainz, 55122 Mainz, Germany
D. Caratelli
University of California Santa Barbara, Santa Barbara, California 93106 USA
D. Carber
Colorado State University, Fort Collins, CO 80523, USA
J. M. Carceller
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
G. Carini
Brookhaven National Laboratory, Upton, NY 11973, USA
B. Carlus
Institut de Physique des 2 Infinis de Lyon, 69622 Villeurbanne, France
M. F. Carneiro
Brookhaven National Laboratory, Upton, NY 11973, USA
P. Carniti
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
I. Caro Terrazas
Colorado State University, Fort Collins, CO 80523, USA
H. Carranza
University of Texas at Arlington, Arlington, TX 76019, USA
N. Carrara
University of California Davis, Davis, CA 95616, USA
L. Carroll
Kansas State University, Manhattan, KS 66506, USA
T. Carroll
University of Wisconsin Madison, Madison, WI 53706, USA
A. Carter
Royal Holloway College London, London, TW20 0EX, United Kingdom
E. Casarejos
University of Vigo, E- 36310 Vigo Spain
D. Casazza
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
J. F. Castaño Forero
Universidad Antonio Nariño, Bogotá, Colombia
F. A. Castaño
University of Antioquia, Medellín, Colombia
A. Castillo
Universidad Sergio Arboleda, 11022 Bogotá, Colombia
C. Castromonte
Universidad Nacional de Ingeniería, Lima 25, Perú
E. Catano-Mur
William and Mary, Williamsburg, VA 23187, USA
C. Cattadori
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
F. Cavalier
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
F. Cavanna
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
S. Centro
Universtà degli Studi di Padova, I-35131 Padova, Italy
G. Cerati
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
C. Cerna
Laboratoire de Physique des Deux Infinis Bordeaux - IN2P3, F-33175 Gradignan, Bordeaux, France,
A. Cervelli
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
A. Cervera Villanueva
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
K. Chakraborty
Physical Research Laboratory, Ahmedabad 380 009, India
S. Chakraborty
Indian Institute of Technology Kanpur, Uttar Pradesh 208016, India
M. Chalifour
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
A. Chappell
University of Warwick, Coventry CV4 7AL, United Kingdom
N. Charitonidis
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
A. Chatterjee
Physical Research Laboratory, Ahmedabad 380 009, India
H. Chen
Brookhaven National Laboratory, Upton, NY 11973, USA
M. Chen
University of California Irvine, Irvine, CA 92697, USA
W. C. Chen
University of Toronto, Toronto, Ontario M5S 1A1, Canada
Y. Chen
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
Z. Chen-Wishart
Royal Holloway College London, London, TW20 0EX, United Kingdom
D. Cherdack
University of Houston, Houston, TX 77204, USA
C. Chi
Columbia University, New York, NY 10027, USA
F. Chiapponi
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
R. Chirco
Illinois Institute of Technology, Chicago, IL 60616, USA
N. Chitirasreemadam
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
Università di Pisa, I-56127 Pisa, Italy
K. Cho
Korea Institute of Science and Technology Information, Daejeon, 34141, South Korea
S. Choate
Northern Illinois University, DeKalb, IL 60115, USA
D. Chokheli
Georgian Technical University, Tbilisi, Georgia
P. S. Chong
University of Pennsylvania, Philadelphia, PA 19104, USA
B. Chowdhury
Argonne National Laboratory, Argonne, IL 60439, USA
D. Christian
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
[…]
M. Chung
Ulsan National Institute of Science and Technology, Ulsan 689-798, South Korea
E. Church
Pacific Northwest National Laboratory, Richland, WA 99352, USA
M. F. Cicala
University College London, London, WC1E 6BT, United Kingdom
M. Cicerchia
Universtà degli Studi di Padova, I-35131 Padova, Italy
V. Cicero
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
R. Ciolini
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
P. Clarke
University of Edinburgh, Edinburgh EH8 9YL, United Kingdom
G. Cline
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
T. E. Coan
Southern Methodist University, Dallas, TX 75275, USA
A. G. Cocco
Istituto Nazionale di Fisica Nucleare Sezione di Napoli, I-80126 Napoli, Italy
J. A. B. Coelho
Université Paris Cité, CNRS, Astroparticule et Cosmologie, Paris, France
A. Cohen
Université Paris Cité, CNRS, Astroparticule et Cosmologie, Paris, France
J. Collazo
University of Vigo, E- 36310 Vigo Spain
J. Collot
University Grenoble Alpes, CNRS, Grenoble INP, LPSC-IN2P3, 38000 Grenoble, France
E. Conley
Duke University, Durham, NC 27708, USA
J. M. Conrad
Massachusetts Institute of Technology, Cambridge, MA 02139, USA
M. Convery
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
S. Copello
Istituto Nazionale di Fisica Nucleare Sezione di Genova, 16146 Genova GE, Italy
P. Cova
Istituto Nazionale di Fisica Nucleare Sezione di Milano, 20133 Milano, Italy
University of Parma, 43121 Parma PR, Italy
C. Cox
Royal Holloway College London, London, TW20 0EX, United Kingdom
L. Cremaldi
University of Mississippi, University, MS 38677 USA
L. Cremonesi
Queen Mary University of London, London E1 4NS, United Kingdom
J. I. Crespo-Anadón
CIEMAT, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas, E-28040 Madrid, Spain
M. Crisler
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
E. Cristaldo
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
Universidad Nacional de Asunción, San Lorenzo, Paraguay
J. Crnkovic
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
G. Crone
University College London, London, WC1E 6BT, United Kingdom
R. Cross
University of Warwick, Coventry CV4 7AL, United Kingdom
A. Cudd
University of Colorado Boulder, Boulder, CO 80309, USA
C. Cuesta
CIEMAT, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas, E-28040 Madrid, Spain
Y. Cui
University of California Riverside, Riverside CA 92521, USA
F. Curciarello
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Frascati, Frascati, Roma, Italy
D. Cussans
University of Bristol, Bristol BS8 1TL, United Kingdom
J. Dai
University Grenoble Alpes, CNRS, Grenoble INP, LPSC-IN2P3, 38000 Grenoble, France
O. Dalager
University of California Irvine, Irvine, CA 92697, USA
R. Dallavalle
Université Paris Cité, CNRS, Astroparticule et Cosmologie, Paris, France
W. Dallaway
University of Toronto, Toronto, Ontario M5S 1A1, Canada
H. da Motta
Centro Brasileiro de Pesquisas Físicas, Rio de Janeiro, RJ 22290-180, Brazil
Z. A. Dar
William and Mary, Williamsburg, VA 23187, USA
R. Darby
University of Sussex, Brighton, BN1 9RH, United Kingdom
L. Da Silva Peres
Universidade Federal do Rio de Janeiro, Rio de Janeiro - RJ, 21941-901, Brazil
Q. David
Institut de Physique des 2 Infinis de Lyon, 69622 Villeurbanne, France
G. S. Davies
University of Mississippi, University, MS 38677 USA
S. Davini
Istituto Nazionale di Fisica Nucleare Sezione di Genova, 16146 Genova GE, Italy
J. Dawson
Université Paris Cité, CNRS, Astroparticule et Cosmologie, Paris, France
R. De Aguiar
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
P. De Almeida
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
P. Debbins
University of Iowa, Iowa City, IA 52242, USA
I. De Bonis
Laboratoire d’Annecy de Physique des Particules, Université Savoie Mont Blanc, CNRS, LAPP-IN2P3, 74000 Annecy, France
M. P. Decowski
Nikhef National Institute of Subatomic Physics, 1098 XG Amsterdam, Netherlands
University of Amsterdam, NL-1098 XG Amsterdam, The Netherlands
A. de Gouvêa
Northwestern University, Evanston, Il 60208, USA
P. C. De Holanda
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
I. L. De Icaza Astiz
University of Sussex, Brighton, BN1 9RH, United Kingdom
P. De Jong
Nikhef National Institute of Subatomic Physics, 1098 XG Amsterdam, Netherlands
University of Amsterdam, NL-1098 XG Amsterdam, The Netherlands
P. Del Amo Sanchez
Laboratoire d’Annecy de Physique des Particules, Université Savoie Mont Blanc, CNRS, LAPP-IN2P3, 74000 Annecy, France
A. De la Torre
CIEMAT, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas, E-28040 Madrid, Spain
G. De Lauretis
Institut de Physique des 2 Infinis de Lyon, 69622 Villeurbanne, France
A. Delbart
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
D. Delepine
Universidad de Guanajuato, Guanajuato, C.P. 37000, Mexico
M. Delgado
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
Università di Milano Bicocca , 20126 Milano, Italy
A. Dell’Acqua
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
G. Delle Monache
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Frascati, Frascati, Roma, Italy
N. Delmonte
Istituto Nazionale di Fisica Nucleare Sezione di Milano, 20133 Milano, Italy
University of Parma, 43121 Parma PR, Italy
P. De Lurgio
Argonne National Laboratory, Argonne, IL 60439, USA
R. Demario
Michigan State University, East Lansing, MI 48824, USA
G. De Matteis
Istituto Nazionale di Fisica Nucleare Sezione di Lecce, 73100 - Lecce, Italy
Università del Salento, 73100 Lecce, Italy
J. R. T. de Mello Neto
Universidade Federal do Rio de Janeiro, Rio de Janeiro - RJ, 21941-901, Brazil
D. M. DeMuth
Valley City State University, Valley City, ND 58072, USA
S. Dennis
University of Cambridge, Cambridge CB3 0HE, United Kingdom
C. Densham
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
P. Denton
Brookhaven National Laboratory, Upton, NY 11973, USA
G. W. Deptuch
Brookhaven National Laboratory, Upton, NY 11973, USA
A. De Roeck
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
V. De Romeri
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
J. P. Detje
University of Cambridge, Cambridge CB3 0HE, United Kingdom
J. Devine
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
R. Dharmapalan
University of Hawaii, Honolulu, HI 96822, USA
M. Dias
Universidade Federal de São Paulo, 09913-030, São Paulo, Brazil
A. Diaz
California Institute of Technology, Pasadena, CA 91125, USA
J. S. Díaz
Indiana University, Bloomington, IN 47405, USA
F. Díaz
Pontificia Universidad Católica del Perú, Lima, Perú
F. Di Capua
Istituto Nazionale di Fisica Nucleare Sezione di Napoli, I-80126 Napoli, Italy
Università degli Studi di Napoli Federico II , 80138 Napoli NA, Italy
A. Di Domenico
Sapienza University of Rome, 00185 Roma RM, Italy
Istituto Nazionale di Fisica Nucleare Sezione di Roma, 00185 Roma RM, Italy
S. Di Domizio
Istituto Nazionale di Fisica Nucleare Sezione di Genova, 16146 Genova GE, Italy
Università degli Studi di Genova, Genova, Italy
S. Di Falco
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
L. Di Giulio
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
P. Ding
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
L. Di Noto
Istituto Nazionale di Fisica Nucleare Sezione di Genova, 16146 Genova GE, Italy
Università degli Studi di Genova, Genova, Italy
E. Diociaiuti
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Frascati, Frascati, Roma, Italy
C. Distefano
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali del Sud, 95123 Catania, Italy
R. Diurba
University of Bern, CH-3012 Bern, Switzerland
M. Diwan
Brookhaven National Laboratory, Upton, NY 11973, USA
Z. Djurcic
Argonne National Laboratory, Argonne, IL 60439, USA
D. Doering
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
S. Dolan
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
F. Dolek
Virginia Tech, Blacksburg, VA 24060, USA
M. J. Dolinski
Drexel University, Philadelphia, PA 19104, USA
D. Domenici
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Frascati, Frascati, Roma, Italy
L. Domine
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
S. Donati
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
Università di Pisa, I-56127 Pisa, Italy
Y. Donon
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
S. Doran
Iowa State University, Ames, Iowa 50011, USA
D. Douglas
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
T.A. Doyle
Stony Brook University, SUNY, Stony Brook, NY 11794, USA
A. Dragone
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
F. Drielsma
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
L. Duarte
Universidade Federal de São Paulo, 09913-030, São Paulo, Brazil
D. Duchesneau
Laboratoire d’Annecy de Physique des Particules, Université Savoie Mont Blanc, CNRS, LAPP-IN2P3, 74000 Annecy, France
K. Duffy
University of Oxford, Oxford, OX1 3RH, United Kingdom
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
K. Dugas
University of California Irvine, Irvine, CA 92697, USA
P. Dunne
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
B. Dutta
Texas A&M University, College Station, Texas 77840
H. Duyang
University of South Carolina, Columbia, SC 29208, USA
D. A. Dwyer
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
A. S. Dyshkant
Northern Illinois University, DeKalb, IL 60115, USA
S. Dytman
University of Pittsburgh, Pittsburgh, PA 15260, USA
M. Eads
Northern Illinois University, DeKalb, IL 60115, USA
A. Earle
University of Sussex, Brighton, BN1 9RH, United Kingdom
S. Edayath
Iowa State University, Ames, Iowa 50011, USA
D. Edmunds
Michigan State University, East Lansing, MI 48824, USA
J. Eisch
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
P. Englezos
Rutgers University, Piscataway, NJ, 08854, USA
A. Ereditato
University of Chicago, Chicago, IL 60637, USA
T. Erjavec
University of California Davis, Davis, CA 95616, USA
C. O. Escobar
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
J. J. Evans
University of Manchester, Manchester M13 9PL, United Kingdom
E. Ewart
Indiana University, Bloomington, IN 47405, USA
A. C. Ezeribe
University of Sheffield, Sheffield S3 7RH, United Kingdom
K. Fahey
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
L. Fajt
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
A. Falcone
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
Università di Milano Bicocca , 20126 Milano, Italy
M. Fani
Los Alamos National Laboratory, Los Alamos, NM 87545, USA
C. Farnese
Istituto Nazionale di Fisica Nucleare Sezione di Padova, 35131 Padova, Italy
S. Farrell
Rice University, Houston, TX 77005
Y. Farzan
Institute for Research in Fundamental Sciences, Tehran, Iran
[…]
J. Felix
Universidad de Guanajuato, Guanajuato, C.P. 37000, Mexico
Y. Feng
Iowa State University, Ames, Iowa 50011, USA
E. Fernandez-Martinez
Madrid Autonoma University and IFT UAM/CSIC, 28049 Madrid, Spain
G. Ferry
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
L. Fields
University of Notre Dame, Notre Dame, IN 46556, USA
P. Filip
Institute of Physics, Czech Academy of Sciences, 182 00 Prague 8, Czech Republic
A. Filkins
Syracuse University, Syracuse, NY 13244, USA
F. Filthaut
Nikhef National Institute of Subatomic Physics, 1098 XG Amsterdam, Netherlands
Radboud University, NL-6525 AJ Nijmegen, Netherlands
R. Fine
Los Alamos National Laboratory, Los Alamos, NM 87545, USA
G. Fiorillo
Istituto Nazionale di Fisica Nucleare Sezione di Napoli, I-80126 Napoli, Italy
Università degli Studi di Napoli Federico II , 80138 Napoli NA, Italy
M. Fiorini
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
University of Ferrara, Ferrara, Italy
S. Fogarty
Colorado State University, Fort Collins, CO 80523, USA
W. Foreman
Illinois Institute of Technology, Chicago, IL 60616, USA
J. Fowler
Duke University, Durham, NC 27708, USA
J. Franc
Czech Technical University, 115 19 Prague 1, Czech Republic
K. Francis
Northern Illinois University, DeKalb, IL 60115, USA
D. Franco
University of Chicago, Chicago, IL 60637, USA
J. Franklin
Durham University, Durham DH1 3LE, United Kingdom
J. Freeman
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
J. Fried
Brookhaven National Laboratory, Upton, NY 11973, USA
A. Friedland
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
S. Fuess
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
I. K. Furic
University of Florida, Gainesville, FL 32611-8440, USA
K. Furman
Queen Mary University of London, London E1 4NS, United Kingdom
A. P. Furmanski
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
R. Gaba
Panjab University, Chandigarh, 160014, India
A. Gabrielli
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
A. M Gago
Pontificia Universidad Católica del Perú, Lima, Perú
F. Galizzi
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
H. Gallagher
Tufts University, Medford, MA 02155, USA
A. Gallas
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
N. Gallice
Brookhaven National Laboratory, Upton, NY 11973, USA
V. Galymov
Institut de Physique des 2 Infinis de Lyon, 69622 Villeurbanne, France
E. Gamberini
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
T. Gamble
University of Sheffield, Sheffield S3 7RH, United Kingdom
F. Ganacim
Universidade Tecnológica Federal do Paraná, Curitiba, Brazil
R. Gandhi
Harish-Chandra Research Institute, Jhunsi, Allahabad 211 019, India
S. Ganguly
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
F. Gao
University of California Santa Barbara, Santa Barbara, California 93106 USA
S. Gao
Brookhaven National Laboratory, Upton, NY 11973, USA
D. Garcia-Gamez
University of Granada & CAFPE, 18002 Granada, Spain
M. Á. García-Peris
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
F. Gardim
Universidade Federal de Alfenas, Poços de Caldas - MG, 37715-400, Brazil
S. Gardiner
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
D. Gastler
Boston University, Boston, MA 02215, USA
A. Gauch
University of Bern, CH-3012 Bern, Switzerland
J. Gauvreau
Occidental College, Los Angeles, CA 90041
P. Gauzzi
Sapienza University of Rome, 00185 Roma RM, Italy
Istituto Nazionale di Fisica Nucleare Sezione di Roma, 00185 Roma RM, Italy
S. Gazzana
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Frascati, Frascati, Roma, Italy
G. Ge
Columbia University, New York, NY 10027, USA
N. Geffroy
Laboratoire d’Annecy de Physique des Particules, Université Savoie Mont Blanc, CNRS, LAPP-IN2P3, 74000 Annecy, France
B. Gelli
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
S. Gent
South Dakota State University, Brookings, SD 57007, USA
L. Gerlach
Brookhaven National Laboratory, Upton, NY 11973, USA
Z. Ghorbani-Moghaddam
Istituto Nazionale di Fisica Nucleare Sezione di Genova, 16146 Genova GE, Italy
T. Giammaria
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
University of Ferrara, Ferrara, Italy
D. Gibin
Universtà degli Studi di Padova, I-35131 Padova, Italy
Istituto Nazionale di Fisica Nucleare Sezione di Padova, 35131 Padova, Italy
I. Gil-Botella
CIEMAT, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas, E-28040 Madrid, Spain
S. Gilligan
Oregon State University, Corvallis, OR 97331, USA
A. Gioiosa
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
S. Giovannella
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Frascati, Frascati, Roma, Italy
C. Girerd
Institut de Physique des 2 Infinis de Lyon, 69622 Villeurbanne, France
A. K. Giri
Indian Institute of Technology Hyderabad, Hyderabad, 502285, India
C. Giugliano
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
V. Giusti
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
D. Gnani
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
O. Gogota
Taras Shevchenko National University of Kyiv, 01601 Kyiv, Ukraine
S. Gollapinni
Los Alamos National Laboratory, Los Alamos, NM 87545, USA
K. Gollwitzer
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
R. A. Gomes
Universidade Federal de Goias, Goiania, GO 74690-900, Brazil
L. V. Gomez Bermeo
Universidad Sergio Arboleda, 11022 Bogotá, Colombia
L. S. Gomez Fajardo
Universidad Sergio Arboleda, 11022 Bogotá, Colombia
F. Gonnella
University of Birmingham, Birmingham B15 2TT, United Kingdom
D. Gonzalez-Diaz
Instituto Galego de Física de Altas Enerxías, University of Santiago de Compostela, Santiago de Compostela, 15782, Spain
M. Gonzalez-Lopez
Madrid Autonoma University and IFT UAM/CSIC, 28049 Madrid, Spain
M. C. Goodman
Argonne National Laboratory, Argonne, IL 60439, USA
S. Goswami
Physical Research Laboratory, Ahmedabad 380 009, India
C. Gotti
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
J. Goudeau
Louisiana State University, Baton Rouge, LA 70803, USA
E. Goudzovski
University of Birmingham, Birmingham B15 2TT, United Kingdom
C. Grace
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
E. Gramellini
University of Manchester, Manchester M13 9PL, United Kingdom
R. Gran
University of Minnesota Duluth, Duluth, MN 55812, USA
E. Granados
Universidad de Guanajuato, Guanajuato, C.P. 37000, Mexico
P. Granger
Université Paris Cité, CNRS, Astroparticule et Cosmologie, Paris, France
C. Grant
Boston University, Boston, MA 02215, USA
D. R. Gratieri
Fluminense Federal University, 9 Icaraí Niterói - RJ, 24220-900, Brazil
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
G. Grauso
Istituto Nazionale di Fisica Nucleare Sezione di Napoli, I-80126 Napoli, Italy
P. Green
University of Oxford, Oxford, OX1 3RH, United Kingdom
S. Greenberg
University of California Berkeley, Berkeley, CA 94720, USA
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
J. Greer
University of Bristol, Bristol BS8 1TL, United Kingdom
W. C. Griffith
University of Sussex, Brighton, BN1 9RH, United Kingdom
F. T. Groetschla
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
K. Grzelak
University of Warsaw, 02-093 Warsaw, Poland
L. Gu
Lancaster University, Lancaster LA1 4YB, United Kingdom
W. Gu
Brookhaven National Laboratory, Upton, NY 11973, USA
V. Guarino
Argonne National Laboratory, Argonne, IL 60439, USA
M. Guarise
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
University of Ferrara, Ferrara, Italy
R. Guenette
University of Manchester, Manchester M13 9PL, United Kingdom
E. Guerard
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
M. Guerzoni
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
D. Guffanti
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
Università di Milano Bicocca , 20126 Milano, Italy
A. Guglielmi
Istituto Nazionale di Fisica Nucleare Sezione di Padova, 35131 Padova, Italy
B. Guo
University of South Carolina, Columbia, SC 29208, USA
Y. Guo
Stony Brook University, SUNY, Stony Brook, NY 11794, USA
A. Gupta
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
V. Gupta
Nikhef National Institute of Subatomic Physics, 1098 XG Amsterdam, Netherlands
University of Amsterdam, NL-1098 XG Amsterdam, The Netherlands
G. Gurung
University of Texas at Arlington, Arlington, TX 76019, USA
D. Gutierrez
University of Puerto Rico, Mayaguez 00681, Puerto Rico, USA
P. Guzowski
University of Manchester, Manchester M13 9PL, United Kingdom
M. M. Guzzo
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
S. Gwon
Chung-Ang University, Seoul 06974, South Korea
A. Habig
University of Minnesota Duluth, Duluth, MN 55812, USA
H. Hadavand
University of Texas at Arlington, Arlington, TX 76019, USA
L. Haegel
Institut de Physique des 2 Infinis de Lyon, 69622 Villeurbanne, France
R. Haenni
University of Bern, CH-3012 Bern, Switzerland
L. Hagaman
Yale University, New Haven, CT 06520, USA
A. Hahn
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
J. Haiston
South Dakota School of Mines and Technology, Rapid City, SD 57701, USA
J. Hakenmüller
Duke University, Durham, NC 27708, USA
T. Hamernik
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
P. Hamilton
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
J. Hancock
University of Birmingham, Birmingham B15 2TT, United Kingdom
F. Happacher
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Frascati, Frascati, Roma, Italy
D. A. Harris
York University, Toronto M3J 1P3, Canada
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
J. Hartnell
University of Sussex, Brighton, BN1 9RH, United Kingdom
T. Hartnett
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
J. Harton
Colorado State University, Fort Collins, CO 80523, USA
T. Hasegawa
High Energy Accelerator Research Organization (KEK), Ibaraki, 305-0801, Japan
C. Hasnip
University of Oxford, Oxford, OX1 3RH, United Kingdom
R. Hatcher
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
K. Hayrapetyan
Queen Mary University of London, London E1 4NS, United Kingdom
J. Hays
Queen Mary University of London, London E1 4NS, United Kingdom
E. Hazen
Boston University, Boston, MA 02215, USA
M. He
University of Houston, Houston, TX 77204, USA
A. Heavey
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
K. M. Heeger
Yale University, New Haven, CT 06520, USA
J. Heise
Sanford Underground Research Facility, Lead, SD, 57754, USA
S. Henry
University of Rochester, Rochester, NY 14627, USA
M. A. Hernandez Morquecho
Illinois Institute of Technology, Chicago, IL 60616, USA
K. Herner
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
V. Hewes
University of Cincinnati, Cincinnati, OH 45221, USA
A. Higuera
Rice University, Houston, TX 77005
C. Hilgenberg
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
S. J. Hillier
University of Birmingham, Birmingham B15 2TT, United Kingdom
A. Himmel
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
E. Hinkle
University of Chicago, Chicago, IL 60637, USA
L.R. Hirsch
Universidade Tecnológica Federal do Paraná, Curitiba, Brazil
J. Ho
Dordt University, Sioux Center, IA 51250, USA
J. Hoff
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Holin
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
T. Holvey
University of Oxford, Oxford, OX1 3RH, United Kingdom
E. Hoppe
Pacific Northwest National Laboratory, Richland, WA 99352, USA
S. Horiuchi
Virginia Tech, Blacksburg, VA 24060, USA
G. A. Horton-Smith
Kansas State University, Manhattan, KS 66506, USA
M. Hostert
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
T. Houdy
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
B. Howard
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
R. Howell
University of Rochester, Rochester, NY 14627, USA
I. Hristova
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
M. S. Hronek
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
J. Huang
University of California Davis, Davis, CA 95616, USA
R.G. Huang
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
Z. Hulcher
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
M. Ibrahim
Eötvös Loránd University, 1053 Budapest, Hungary
G. Iles
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
N. Ilic
University of Toronto, Toronto, Ontario M5S 1A1, Canada
A. M. Iliescu
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Frascati, Frascati, Roma, Italy
R. Illingworth
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
G. Ingratta
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
A. Ioannisian
Yerevan Institute for Theoretical Physics and Modeling, Yerevan 0036, Armenia
B. Irwin
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
L. Isenhower
Abilene Christian University, Abilene, TX 79601, USA
M. Ismerio Oliveira
Universidade Federal do Rio de Janeiro, Rio de Janeiro - RJ, 21941-901, Brazil
R. Itay
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
C.M. Jackson
Pacific Northwest National Laboratory, Richland, WA 99352, USA
V. Jain
University of Albany, SUNY, Albany, NY 12222, USA
E. James
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
W. Jang
University of Texas at Arlington, Arlington, TX 76019, USA
B. Jargowsky
University of California Irvine, Irvine, CA 92697, USA
D. Jena
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
I. Jentz
University of Wisconsin Madison, Madison, WI 53706, USA
X. Ji
Brookhaven National Laboratory, Upton, NY 11973, USA
C. Jiang
Jackson State University, Jackson, MS 39217, USA
J. Jiang
Stony Brook University, SUNY, Stony Brook, NY 11794, USA
L. Jiang
Virginia Tech, Blacksburg, VA 24060, USA
A. Jipa
University of Bucharest, Bucharest, Romania
F. R. Joaquim
Laboratório de Instrumentação e Física Experimental de Partículas, 1649-003 Lisboa and 3004-516 Coimbra, Portugal
Instituto Superior Técnico - IST, Universidade de Lisboa, Portugal
W. Johnson
South Dakota School of Mines and Technology, Rapid City, SD 57701, USA
C. Jollet
Laboratoire de Physique des Deux Infinis Bordeaux - IN2P3, F-33175 Gradignan, Bordeaux, France,
B. Jones
University of Texas at Arlington, Arlington, TX 76019, USA
R. Jones
University of Sheffield, Sheffield S3 7RH, United Kingdom
D. José Fernández
Instituto Galego de Física de Altas Enerxías, University of Santiago de Compostela, Santiago de Compostela, 15782, Spain
N. Jovancevic
University of Novi Sad, 21102 Novi Sad, Serbia
M. Judah
University of Pittsburgh, Pittsburgh, PA 15260, USA
C. K. Jung
Stony Brook University, SUNY, Stony Brook, NY 11794, USA
T. Junk
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
Y. Jwa
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
Columbia University, New York, NY 10027, USA
M. Kabirnezhad
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
A. C. Kaboth
Royal Holloway College London, London, TW20 0EX, United Kingdom
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
I. Kadenko
Taras Shevchenko National University of Kyiv, 01601 Kyiv, Ukraine
[…]
D. Kalra
Columbia University, New York, NY 10027, USA
M. Kandemir
Erciyes University, Kayseri, Turkey
D. M. Kaplan
Illinois Institute of Technology, Chicago, IL 60616, USA
G. Karagiorgi
Columbia University, New York, NY 10027, USA
G. Karaman
University of Iowa, Iowa City, IA 52242, USA
A. Karcher
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
Y. Karyotakis
Laboratoire d’Annecy de Physique des Particules, Université Savoie Mont Blanc, CNRS, LAPP-IN2P3, 74000 Annecy, France
S. Kasai
National Institute of Technology, Kure College, Hiroshima, 737-8506, Japan
S. P. Kasetti
Louisiana State University, Baton Rouge, LA 70803, USA
L. Kashur
Colorado State University, Fort Collins, CO 80523, USA
I. Katsioulas
University of Birmingham, Birmingham B15 2TT, United Kingdom
A. Kauther
Northern Illinois University, DeKalb, IL 60115, USA
N. Kazaryan
Yerevan Institute for Theoretical Physics and Modeling, Yerevan 0036, Armenia
L. Ke
Brookhaven National Laboratory, Upton, NY 11973, USA
E. Kearns
Boston University, Boston, MA 02215, USA
P.T. Keener
University of Pennsylvania, Philadelphia, PA 19104, USA
K.J. Kelly
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
E. Kemp
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
O. Kemularia
Georgian Technical University, Tbilisi, Georgia
Y. Kermaidic
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
W. Ketchum
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
S. H. Kettell
Brookhaven National Laboratory, Upton, NY 11973, USA
[…]
N. Khan
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
A. Khvedelidze
Georgian Technical University, Tbilisi, Georgia
D. Kim
Texas A&M University, College Station, Texas 77840
J. Kim
University of Rochester, Rochester, NY 14627, USA
B. King
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
B. Kirby
Columbia University, New York, NY 10027, USA
M. Kirby
Brookhaven National Laboratory, Upton, NY 11973, USA
A. Kish
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
J. Klein
University of Pennsylvania, Philadelphia, PA 19104, USA
J. Kleykamp
University of Mississippi, University, MS 38677 USA
A. Klustova
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
T. Kobilarcik
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
L. Koch
Johannes Gutenberg-Universität Mainz, 55122 Mainz, Germany
K. Koehler
University of Wisconsin Madison, Madison, WI 53706, USA
L. W. Koerner
University of Houston, Houston, TX 77204, USA
D. H. Koh
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
[…]
M. Kordosky
William and Mary, Williamsburg, VA 23187, USA
T. Kosc
University Grenoble Alpes, CNRS, Grenoble INP, LPSC-IN2P3, 38000 Grenoble, France
U. Kose
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
V. A. Kostelecký
Indiana University, Bloomington, IN 47405, USA
K. Kothekar
University of Bristol, Bristol BS8 1TL, United Kingdom
I. Kotler
Drexel University, Philadelphia, PA 19104, USA
M. Kovalcuk
Institute of Physics, Czech Academy of Sciences, 182 00 Prague 8, Czech Republic
[…]
W. Krah
Nikhef National Institute of Subatomic Physics, 1098 XG Amsterdam, Netherlands
R. Kralik
University of Sussex, Brighton, BN1 9RH, United Kingdom
M. Kramer
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
L. Kreczko
University of Bristol, Bristol BS8 1TL, United Kingdom
F. Krennrich
Iowa State University, Ames, Iowa 50011, USA
I. Kreslo
University of Bern, CH-3012 Bern, Switzerland
T. Kroupova
University of Pennsylvania, Philadelphia, PA 19104, USA
S. Kubota
University of Manchester, Manchester M13 9PL, United Kingdom
M. Kubu
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
[…]
V. A. Kudryavtsev
University of Sheffield, Sheffield S3 7RH, United Kingdom
G. Kufatty
Florida State University, Tallahassee, FL, 32306 USA
S. Kuhlmann
Argonne National Laboratory, Argonne, IL 60439, USA
J. Kumar
University of Hawaii, Honolulu, HI 96822, USA
P. Kumar
University of Sheffield, Sheffield S3 7RH, United Kingdom
S. Kumaran
University of California Irvine, Irvine, CA 92697, USA
P. Kunze
Laboratoire d’Annecy de Physique des Particules, Université Savoie Mont Blanc, CNRS, LAPP-IN2P3, 74000 Annecy, France
J. Kunzmann
University of Bern, CH-3012 Bern, Switzerland
R. Kuravi
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
N. Kurita
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
C. Kuruppu
University of South Carolina, Columbia, SC 29208, USA
V. Kus
Czech Technical University, 115 19 Prague 1, Czech Republic
T. Kutter
Louisiana State University, Baton Rouge, LA 70803, USA
J. Kvasnicka
Institute of Physics, Czech Academy of Sciences, 182 00 Prague 8, Czech Republic
T. Labree
Northern Illinois University, DeKalb, IL 60115, USA
T. Lackey
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Lambert
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
B. J. Land
University of Pennsylvania, Philadelphia, PA 19104, USA
C. E. Lane
Drexel University, Philadelphia, PA 19104, USA
N. Lane
University of Manchester, Manchester M13 9PL, United Kingdom
K. Lang
University of Texas at Austin, Austin, TX 78712, USA
T. Langford
Yale University, New Haven, CT 06520, USA
M. Langstaff
University of Manchester, Manchester M13 9PL, United Kingdom
F. Lanni
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
O. Lantwin
Laboratoire d’Annecy de Physique des Particules, Université Savoie Mont Blanc, CNRS, LAPP-IN2P3, 74000 Annecy, France
J. Larkin
Brookhaven National Laboratory, Upton, NY 11973, USA
P. Lasorak
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
D. Last
University of Pennsylvania, Philadelphia, PA 19104, USA
A. Laudrain
Johannes Gutenberg-Universität Mainz, 55122 Mainz, Germany
A. Laundrie
University of Wisconsin Madison, Madison, WI 53706, USA
G. Laurenti
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
E. Lavaut
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
A. Lawrence
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
P. Laycock
Brookhaven National Laboratory, Upton, NY 11973, USA
I. Lazanu
University of Bucharest, Bucharest, Romania
M. Lazzaroni
Istituto Nazionale di Fisica Nucleare Sezione di Milano, 20133 Milano, Italy
Università degli Studi di Milano, I-20133 Milano, Italy
T. Le
Tufts University, Medford, MA 02155, USA
S. Leardini
Instituto Galego de Física de Altas Enerxías, University of Santiago de Compostela, Santiago de Compostela, 15782, Spain
J. Learned
University of Hawaii, Honolulu, HI 96822, USA
T. LeCompte
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
C. Lee
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
V. Legin
Taras Shevchenko National University of Kyiv, 01601 Kyiv, Ukraine
G. Lehmann Miotto
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
R. Lehnert
Indiana University, Bloomington, IN 47405, USA
M. A. Leigui de Oliveira
Universidade Federal do ABC, Santo André - SP, 09210-580, Brazil
M. Leitner
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
D. Leon Silverio
South Dakota School of Mines and Technology, Rapid City, SD 57701, USA
L. M. Lepin
Florida State University, Tallahassee, FL, 32306 USA
University of Manchester, Manchester M13 9PL, United Kingdom
J.-Y Li
University of Edinburgh, Edinburgh EH8 9YL, United Kingdom
S. W. Li
University of California Davis, Davis, CA 95616, USA
Y. Li
Brookhaven National Laboratory, Upton, NY 11973, USA
H. Liao
Kansas State University, Manhattan, KS 66506, USA
C. S. Lin
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
D. Lindebaum
University of Bristol, Bristol BS8 1TL, United Kingdom
S. Linden
Brookhaven National Laboratory, Upton, NY 11973, USA
R. A. Lineros
Universidad Católica del Norte, Antofagasta, Chile
J. Ling
Sun Yat-Sen University, Guangzhou, 510275, China
A. Lister
University of Wisconsin Madison, Madison, WI 53706, USA
B. R. Littlejohn
Illinois Institute of Technology, Chicago, IL 60616, USA
H. Liu
Brookhaven National Laboratory, Upton, NY 11973, USA
J. Liu
University of California Irvine, Irvine, CA 92697, USA
Y. Liu
University of Chicago, Chicago, IL 60637, USA
S. Lockwitz
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
M. Lokajicek
Institute of Physics, Czech Academy of Sciences, 182 00 Prague 8, Czech Republic
I. Lomidze
Georgian Technical University, Tbilisi, Georgia
K. Long
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
T. V. Lopes
Universidade Federal de Alfenas, Poços de Caldas - MG, 37715-400, Brazil
J.Lopez
University of Antioquia, Medellín, Colombia
I. López de Rego
CIEMAT, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas, E-28040 Madrid, Spain
N. López-March
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
T. Lord
University of Warwick, Coventry CV4 7AL, United Kingdom
J. M. LoSecco
University of Notre Dame, Notre Dame, IN 46556, USA
W. C. Louis
Los Alamos National Laboratory, Los Alamos, NM 87545, USA
A. Lozano Sanchez
Drexel University, Philadelphia, PA 19104, USA
X.-G. Lu
University of Warwick, Coventry CV4 7AL, United Kingdom
K.B. Luk
Hong Kong University of Science and Technology, Kowloon, Hong Kong, China
University of California Berkeley, Berkeley, CA 94720, USA
B. Lunday
University of Pennsylvania, Philadelphia, PA 19104, USA
X. Luo
University of California Santa Barbara, Santa Barbara, California 93106 USA
E. Luppi
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
University of Ferrara, Ferrara, Italy
J. Maalmi
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
D. MacFarlane
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
A. A. Machado
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
P. Machado
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
C. T. Macias
Indiana University, Bloomington, IN 47405, USA
J. R. Macier
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
M. MacMahon
University College London, London, WC1E 6BT, United Kingdom
A. Maddalena
Laboratori Nazionali del Gran Sasso, L’Aquila AQ, Italy
A. Madera
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
P. Madigan
University of California Berkeley, Berkeley, CA 94720, USA
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
S. Magill
Argonne National Laboratory, Argonne, IL 60439, USA
C. Magueur
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
K. Mahn
Michigan State University, East Lansing, MI 48824, USA
A. Maio
Laboratório de Instrumentação e Física Experimental de Partículas, 1649-003 Lisboa and 3004-516 Coimbra, Portugal
Faculdade de Ciências da Universidade de Lisboa - FCUL, 1749-016 Lisboa, Portugal
A. Major
Duke University, Durham, NC 27708, USA
K. Majumdar
University of Liverpool, L69 7ZE, Liverpool, United Kingdom
M. Man
University of Toronto, Toronto, Ontario M5S 1A1, Canada
R. C. Mandujano
University of California Irvine, Irvine, CA 92697, USA
J. Maneira
Laboratório de Instrumentação e Física Experimental de Partículas, 1649-003 Lisboa and 3004-516 Coimbra, Portugal
Faculdade de Ciências da Universidade de Lisboa - FCUL, 1749-016 Lisboa, Portugal
S. Manly
University of Rochester, Rochester, NY 14627, USA
A. Mann
Tufts University, Medford, MA 02155, USA
K. Manolopoulos
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
M. Manrique Plata
Indiana University, Bloomington, IN 47405, USA
S. Manthey Corchado
CIEMAT, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas, E-28040 Madrid, Spain
V. N. Manyam
Brookhaven National Laboratory, Upton, NY 11973, USA
M. Marchan
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Marchionni
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
W. Marciano
Brookhaven National Laboratory, Upton, NY 11973, USA
D. Marfatia
University of Hawaii, Honolulu, HI 96822, USA
C. Mariani
Virginia Tech, Blacksburg, VA 24060, USA
J. Maricic
University of Hawaii, Honolulu, HI 96822, USA
F. Marinho
Instituto Tecnológico de Aeronáutica, Sao Jose dos Campos, Brazil
A. D. Marino
University of Colorado Boulder, Boulder, CO 80309, USA
T. Markiewicz
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
F. Das Chagas Marques
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
C. Marquet
Laboratoire de Physique des Deux Infinis Bordeaux - IN2P3, F-33175 Gradignan, Bordeaux, France,
D. Marsden
University of Manchester, Manchester M13 9PL, United Kingdom
M. Marshak
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
C. M. Marshall
University of Rochester, Rochester, NY 14627, USA
J. Marshall
University of Warwick, Coventry CV4 7AL, United Kingdom
L. Martina
Istituto Nazionale di Fisica Nucleare Sezione di Lecce, 73100 - Lecce, Italy
Università del Salento, 73100 Lecce, Italy
J. Martín-Albo
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
N. Martinez
Kansas State University, Manhattan, KS 66506, USA
D.A. Martinez Caicedo
South Dakota School of Mines and Technology, Rapid City, SD 57701, USA
F. Martínez López
Queen Mary University of London, London E1 4NS, United Kingdom
P. Martínez Miravé
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
S. Martynenko
Brookhaven National Laboratory, Upton, NY 11973, USA
V. Mascagna
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
C. Massari
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
A. Mastbaum
Rutgers University, Piscataway, NJ, 08854, USA
F. Matichard
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
S. Matsuno
University of Hawaii, Honolulu, HI 96822, USA
G. Matteucci
Istituto Nazionale di Fisica Nucleare Sezione di Napoli, I-80126 Napoli, Italy
Università degli Studi di Napoli Federico II , 80138 Napoli NA, Italy
J. Matthews
Louisiana State University, Baton Rouge, LA 70803, USA
C. Mauger
University of Pennsylvania, Philadelphia, PA 19104, USA
N. Mauri
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
K. Mavrokoridis
University of Liverpool, L69 7ZE, Liverpool, United Kingdom
I. Mawby
Lancaster University, Lancaster LA1 4YB, United Kingdom
R. Mazza
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
A. Mazzacane
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
T. McAskill
Wellesley College, Wellesley, MA 02481, USA
N. McConkey
University College London, London, WC1E 6BT, United Kingdom
K. S. McFarland
University of Rochester, Rochester, NY 14627, USA
C. McGrew
Stony Brook University, SUNY, Stony Brook, NY 11794, USA
A. McNab
University of Manchester, Manchester M13 9PL, United Kingdom
L. Meazza
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
V. C. N. Meddage
University of Florida, Gainesville, FL 32611-8440, USA
B. Mehta
Panjab University, Chandigarh, 160014, India
P. Mehta
Jawaharlal Nehru University, New Delhi 110067, India
P. Melas
University of Athens, Zografou GR 157 84, Greece
O. Mena
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
H. Mendez
University of Puerto Rico, Mayaguez 00681, Puerto Rico, USA
P. Mendez
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
D. P. Méndez
Brookhaven National Laboratory, Upton, NY 11973, USA
A. Menegolli
Istituto Nazionale di Fisica Nucleare Sezione di Pavia, I-27100 Pavia, Italy
Università degli Studi di Pavia, 27100 Pavia PV, Italy
G. Meng
Istituto Nazionale di Fisica Nucleare Sezione di Padova, 35131 Padova, Italy
A. C. E. A. Mercuri
Universidade Tecnológica Federal do Paraná, Curitiba, Brazil
A. Meregaglia
Laboratoire de Physique des Deux Infinis Bordeaux - IN2P3, F-33175 Gradignan, Bordeaux, France,
M. D. Messier
Indiana University, Bloomington, IN 47405, USA
S. Metallo
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
J. Metcalf
Tufts University, Medford, MA 02155, USA
Massachusetts Institute of Technology, Cambridge, MA 02139, USA
W. Metcalf
Louisiana State University, Baton Rouge, LA 70803, USA
M. Mewes
Indiana University, Bloomington, IN 47405, USA
H. Meyer
Wichita State University, Wichita, KS 67260, USA
T. Miao
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Miccoli
Istituto Nazionale di Fisica Nucleare Sezione di Lecce, 73100 - Lecce, Italy
G. Michna
South Dakota State University, Brookings, SD 57007, USA
V. Mikola
University College London, London, WC1E 6BT, United Kingdom
R. Milincic
University of Hawaii, Honolulu, HI 96822, USA
F. Miller
University of Wisconsin Madison, Madison, WI 53706, USA
G. Miller
University of Manchester, Manchester M13 9PL, United Kingdom
W. Miller
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
[…]
A. Minotti
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
Università di Milano Bicocca , 20126 Milano, Italy
L. Miralles
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
O. G. Miranda
Centro de Investigación y de Estudios Avanzados del Instituto Politécnico Nacional (Cinvestav), Mexico City, Mexico
C. Mironov
Université Paris Cité, CNRS, Astroparticule et Cosmologie, Paris, France
S. Miryala
Brookhaven National Laboratory, Upton, NY 11973, USA
S. Miscetti
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Frascati, Frascati, Roma, Italy
C. S. Mishra
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
S. R. Mishra
University of South Carolina, Columbia, SC 29208, USA
A. Mislivec
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
M. Mitchell
Louisiana State University, Baton Rouge, LA 70803, USA
D. Mladenov
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
I. Mocioiu
Pennsylvania State University, University Park, PA 16802, USA
A. Mogan
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
N. Moggi
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
R. Mohanta
University of Hyderabad, Gachibowli, Hyderabad - 500 046, India
T. A. Mohayai
Indiana University, Bloomington, IN 47405, USA
N. Mokhov
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
J. Molina
Universidad Nacional de Asunción, San Lorenzo, Paraguay
L. Molina Bueno
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
E. Montagna
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
A. Montanari
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
C. Montanari
Istituto Nazionale di Fisica Nucleare Sezione di Pavia, I-27100 Pavia, Italy
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
Università degli Studi di Pavia, 27100 Pavia PV, Italy
D. Montanari
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
D. Montanino
Istituto Nazionale di Fisica Nucleare Sezione di Lecce, 73100 - Lecce, Italy
Università del Salento, 73100 Lecce, Italy
L. M. Montaño Zetina
Centro de Investigación y de Estudios Avanzados del Instituto Politécnico Nacional (Cinvestav), Mexico City, Mexico
M. Mooney
Colorado State University, Fort Collins, CO 80523, USA
A. F. Moor
University of Sheffield, Sheffield S3 7RH, United Kingdom
Z. Moore
Syracuse University, Syracuse, NY 13244, USA
D. Moreno
Universidad Antonio Nariño, Bogotá, Colombia
O. Moreno-Palacios
William and Mary, Williamsburg, VA 23187, USA
L. Morescalchi
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
D. Moretti
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
R. Moretti
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
C. Morris
University of Houston, Houston, TX 77204, USA
C. Mossey
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
M. Mote
Louisiana State University, Baton Rouge, LA 70803, USA
C. A. Moura
Universidade Federal do ABC, Santo André - SP, 09210-580, Brazil
G. Mouster
Lancaster University, Lancaster LA1 4YB, United Kingdom
W. Mu
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
L. Mualem
California Institute of Technology, Pasadena, CA 91125, USA
J. Mueller
Colorado State University, Fort Collins, CO 80523, USA
M. Muether
Wichita State University, Wichita, KS 67260, USA
F. Muheim
University of Edinburgh, Edinburgh EH8 9YL, United Kingdom
A. Muir
Daresbury Laboratory, Cheshire WA4 4AD, United Kingdom
M. Mulhearn
University of California Davis, Davis, CA 95616, USA
D. Munford
University of Houston, Houston, TX 77204, USA
L. J. Munteanu
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
H. Muramatsu
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
J. Muraz
University Grenoble Alpes, CNRS, Grenoble INP, LPSC-IN2P3, 38000 Grenoble, France
M. Murphy
Virginia Tech, Blacksburg, VA 24060, USA
T. Murphy
Syracuse University, Syracuse, NY 13244, USA
J. Muse
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
A. Mytilinaki
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
J. Nachtman
University of Iowa, Iowa City, IA 52242, USA
Y. Nagai
Eötvös Loránd University, 1053 Budapest, Hungary
S. Nagu
University of Lucknow, Uttar Pradesh 226007, India
R. Nandakumar
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
D. Naples
University of Pittsburgh, Pittsburgh, PA 15260, USA
S. Narita
Iwate University, Morioka, Iwate 020-8551, Japan
A. Nath
Indian Institute of Technology Guwahati, Guwahati, 781 039, India
A. Navrer-Agasson
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
N. Nayak
Brookhaven National Laboratory, Upton, NY 11973, USA
M. Nebot-Guinot
University of Edinburgh, Edinburgh EH8 9YL, United Kingdom
A. Nehm
Johannes Gutenberg-Universität Mainz, 55122 Mainz, Germany
J. K. Nelson
William and Mary, Williamsburg, VA 23187, USA
O. Neogi
University of Iowa, Iowa City, IA 52242, USA
J. Nesbit
University of Wisconsin Madison, Madison, WI 53706, USA
M. Nessi
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
D. Newbold
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
M. Newcomer
University of Pennsylvania, Philadelphia, PA 19104, USA
R. Nichol
University College London, London, WC1E 6BT, United Kingdom
F. Nicolas-Arnaldos
University of Granada & CAFPE, 18002 Granada, Spain
A. Nikolica
University of Pennsylvania, Philadelphia, PA 19104, USA
J. Nikolov
University of Novi Sad, 21102 Novi Sad, Serbia
E. Niner
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
K. Nishimura
University of Hawaii, Honolulu, HI 96822, USA
A. Norman
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Norrick
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
P. Novella
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
J. A. Nowak
Lancaster University, Lancaster LA1 4YB, United Kingdom
M. Oberling
Argonne National Laboratory, Argonne, IL 60439, USA
J. P. Ochoa-Ricoux
University of California Irvine, Irvine, CA 92697, USA
S. Oh
Duke University, Durham, NC 27708, USA
S.B. Oh
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Olivier
University of Notre Dame, Notre Dame, IN 46556, USA
[…]
T. Olson
University of Houston, Houston, TX 77204, USA
Y. Onel
University of Iowa, Iowa City, IA 52242, USA
Y. Onishchuk
Taras Shevchenko National University of Kyiv, 01601 Kyiv, Ukraine
A. Oranday
Indiana University, Bloomington, IN 47405, USA
M. Osbiston
University of Warwick, Coventry CV4 7AL, United Kingdom
J. A. Osorio Vélez
University of Antioquia, Medellín, Colombia
L. Otiniano Ormachea
Comisión Nacional de Investigación y Desarrollo Aeroespacial, Lima, Peru
Universidad Nacional de Ingeniería, Lima 25, Perú
J. Ott
University of California Irvine, Irvine, CA 92697, USA
L. Pagani
University of California Davis, Davis, CA 95616, USA
G. Palacio
Universidad EIA, Envigado, Antioquia, Colombia
O. Palamara
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
S. Palestini
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
J. M. Paley
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
M. Pallavicini
Istituto Nazionale di Fisica Nucleare Sezione di Genova, 16146 Genova GE, Italy
Università degli Studi di Genova, Genova, Italy
C. Palomares
CIEMAT, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas, E-28040 Madrid, Spain
S. Pan
Physical Research Laboratory, Ahmedabad 380 009, India
P. Panda
University of Hyderabad, Gachibowli, Hyderabad - 500 046, India
W. Panduro Vazquez
Royal Holloway College London, London, TW20 0EX, United Kingdom
E. Pantic
University of California Davis, Davis, CA 95616, USA
V. Paolone
University of Pittsburgh, Pittsburgh, PA 15260, USA
V. Papadimitriou
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
R. Papaleo
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali del Sud, 95123 Catania, Italy
A. Papanestis
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
D. Papoulias
University of Athens, Zografou GR 157 84, Greece
S. Paramesvaran
University of Bristol, Bristol BS8 1TL, United Kingdom
A. Paris
University of Puerto Rico, Mayaguez 00681, Puerto Rico, USA
S. Parke
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
E. Parozzi
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
Università di Milano Bicocca , 20126 Milano, Italy
S. Parsa
University of Bern, CH-3012 Bern, Switzerland
Z. Parsa
Brookhaven National Laboratory, Upton, NY 11973, USA
S. Parveen
Jawaharlal Nehru University, New Delhi 110067, India
M. Parvu
University of Bucharest, Bucharest, Romania
D. Pasciuto
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
S. Pascoli
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
L. Pasqualini
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
J. Pasternak
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
C. Patrick
University of Edinburgh, Edinburgh EH8 9YL, United Kingdom
University College London, London, WC1E 6BT, United Kingdom
L. Patrizii
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
R. B. Patterson
California Institute of Technology, Pasadena, CA 91125, USA
T. Patzak
Université Paris Cité, CNRS, Astroparticule et Cosmologie, Paris, France
A. Paudel
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
L. Paulucci
Universidade Federal do ABC, Santo André - SP, 09210-580, Brazil
Z. Pavlovic
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
G. Pawloski
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
D. Payne
University of Liverpool, L69 7ZE, Liverpool, United Kingdom
V. Pec
Institute of Physics, Czech Academy of Sciences, 182 00 Prague 8, Czech Republic
E. Pedreschi
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
S. J. M. Peeters
University of Sussex, Brighton, BN1 9RH, United Kingdom
W. Pellico
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Pena Perez
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
E. Pennacchio
Institut de Physique des 2 Infinis de Lyon, 69622 Villeurbanne, France
A. Penzo
University of Iowa, Iowa City, IA 52242, USA
O. L. G. Peres
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
Y. F. Perez Gonzalez
Durham University, Durham DH1 3LE, United Kingdom
L. Pérez-Molina
CIEMAT, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas, E-28040 Madrid, Spain
C. Pernas
William and Mary, Williamsburg, VA 23187, USA
J. Perry
University of Edinburgh, Edinburgh EH8 9YL, United Kingdom
D. Pershey
Florida State University, Tallahassee, FL, 32306 USA
G. Pessina
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
G. Petrillo
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
C. Petta
Istituto Nazionale di Fisica Nucleare Sezione di Catania, I-95123 Catania, Italy
Università di Catania, 2 - 95131 Catania, Italy
R. Petti
University of South Carolina, Columbia, SC 29208, USA
M. Pfaff
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
V. Pia
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
L. Pickering
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
Royal Holloway College London, London, TW20 0EX, United Kingdom
F. Pietropaolo
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
Istituto Nazionale di Fisica Nucleare Sezione di Padova, 35131 Padova, Italy
V.L.Pimentel
Centro de Tecnologia da Informacao Renato Archer, Amarais - Campinas, SP - CEP 13069-901
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
G. Pinaroli
Brookhaven National Laboratory, Upton, NY 11973, USA
J. Pinchault
Laboratoire d’Annecy de Physique des Particules, Université Savoie Mont Blanc, CNRS, LAPP-IN2P3, 74000 Annecy, France
K. Pitts
Virginia Tech, Blacksburg, VA 24060, USA
K. Plows
University of Oxford, Oxford, OX1 3RH, United Kingdom
R. Plunkett
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
C. Pollack
University of Puerto Rico, Mayaguez 00681, Puerto Rico, USA
T. Pollman
Nikhef National Institute of Subatomic Physics, 1098 XG Amsterdam, Netherlands
University of Amsterdam, NL-1098 XG Amsterdam, The Netherlands
D. Polo-Toledo
Universidad del Atlántico, Barranquilla, Atlántico, Colombia
F. Pompa
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
X. Pons
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
N. Poonthottathil
Indian Institute of Technology Kanpur, Uttar Pradesh 208016, India
Iowa State University, Ames, Iowa 50011, USA
V. Popov
Tel Aviv University, Tel Aviv-Yafo, Israel
F. Poppi
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
J. Porter
University of Sussex, Brighton, BN1 9RH, United Kingdom
M. Potekhin
Brookhaven National Laboratory, Upton, NY 11973, USA
R. Potenza
Istituto Nazionale di Fisica Nucleare Sezione di Catania, I-95123 Catania, Italy
Università di Catania, 2 - 95131 Catania, Italy
J. Pozimski
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
M. Pozzato
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
T. Prakash
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
C. Pratt
University of California Davis, Davis, CA 95616, USA
M. Prest
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
F. Psihas
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
D. Pugnere
Institut de Physique des 2 Infinis de Lyon, 69622 Villeurbanne, France
X. Qian
Brookhaven National Laboratory, Upton, NY 11973, USA
J. Queen
Duke University, Durham, NC 27708, USA
J. L. Raaf
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
V. Radeka
Brookhaven National Laboratory, Upton, NY 11973, USA
J. Rademacker
University of Bristol, Bristol BS8 1TL, United Kingdom
B. Radics
York University, Toronto M3J 1P3, Canada
A. Rafique
Argonne National Laboratory, Argonne, IL 60439, USA
E. Raguzin
Brookhaven National Laboratory, Upton, NY 11973, USA
M. Rai
University of Warwick, Coventry CV4 7AL, United Kingdom
S. Rajagopalan
Brookhaven National Laboratory, Upton, NY 11973, USA
M. Rajaoalisoa
University of Cincinnati, Cincinnati, OH 45221, USA
I. Rakhno
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
L. Rakotondravohitra
University of Antananarivo, Antananarivo 101, Madagascar
L. Ralte
Indian Institute of Technology Hyderabad, Hyderabad, 502285, India
M. A. Ramirez Delgado
University of Pennsylvania, Philadelphia, PA 19104, USA
B. Ramson
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Rappoldi
Istituto Nazionale di Fisica Nucleare Sezione di Pavia, I-27100 Pavia, Italy
Università degli Studi di Pavia, 27100 Pavia PV, Italy
G. Raselli
Istituto Nazionale di Fisica Nucleare Sezione di Pavia, I-27100 Pavia, Italy
Università degli Studi di Pavia, 27100 Pavia PV, Italy
P. Ratoff
Lancaster University, Lancaster LA1 4YB, United Kingdom
R. Ray
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
H. Razafinime
University of Cincinnati, Cincinnati, OH 45221, USA
E. M. Rea
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
J. S. Real
University Grenoble Alpes, CNRS, Grenoble INP, LPSC-IN2P3, 38000 Grenoble, France
B. Rebel
University of Wisconsin Madison, Madison, WI 53706, USA
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
R. Rechenmacher
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
M. Reggiani-Guzzo
University of Manchester, Manchester M13 9PL, United Kingdom
J. Reichenbacher
South Dakota School of Mines and Technology, Rapid City, SD 57701, USA
S. D. Reitzner
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
H. Rejeb Sfar
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
E. Renner
Los Alamos National Laboratory, Los Alamos, NM 87545, USA
A. Renshaw
University of Houston, Houston, TX 77204, USA
S. Rescia
Brookhaven National Laboratory, Upton, NY 11973, USA
F. Resnati
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
Diego Restrepo
University of Antioquia, Medellín, Colombia
C. Reynolds
Queen Mary University of London, London E1 4NS, United Kingdom
M. Ribas
Universidade Tecnológica Federal do Paraná, Curitiba, Brazil
S. Riboldi
Istituto Nazionale di Fisica Nucleare Sezione di Milano, 20133 Milano, Italy
C. Riccio
Stony Brook University, SUNY, Stony Brook, NY 11794, USA
G. Riccobene
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali del Sud, 95123 Catania, Italy
J. S. Ricol
University Grenoble Alpes, CNRS, Grenoble INP, LPSC-IN2P3, 38000 Grenoble, France
M. Rigan
University of Sussex, Brighton, BN1 9RH, United Kingdom
E. V. Rincón
Universidad EIA, Envigado, Antioquia, Colombia
A. Ritchie-Yates
Royal Holloway College London, London, TW20 0EX, United Kingdom
S. Ritter
Johannes Gutenberg-Universität Mainz, 55122 Mainz, Germany
D. Rivera
Los Alamos National Laboratory, Los Alamos, NM 87545, USA
R. Rivera
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Robert
University Grenoble Alpes, CNRS, Grenoble INP, LPSC-IN2P3, 38000 Grenoble, France
J. L. Rocabado Rocha
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
L. Rochester
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
M. Roda
University of Liverpool, L69 7ZE, Liverpool, United Kingdom
P. Rodrigues
University of Oxford, Oxford, OX1 3RH, United Kingdom
M. J. Rodriguez Alonso
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
J. Rodriguez Rondon
South Dakota School of Mines and Technology, Rapid City, SD 57701, USA
A. J. Roeth
Duke University, Durham, NC 27708, USA
S. Rosauro-Alcaraz
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
P. Rosier
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
D. Ross
Michigan State University, East Lansing, MI 48824, USA
M. Rossella
Istituto Nazionale di Fisica Nucleare Sezione di Pavia, I-27100 Pavia, Italy
Università degli Studi di Pavia, 27100 Pavia PV, Italy
M. Rossi
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
M. Ross-Lonergan
Los Alamos National Laboratory, Los Alamos, NM 87545, USA
N. Roy
York University, Toronto M3J 1P3, Canada
P. Roy
Wichita State University, Wichita, KS 67260, USA
C. Rubbia
Gran Sasso Science Institute, L’Aquila, Italy
A. Ruggeri
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
G. Ruiz Ferreira
University of Manchester, Manchester M13 9PL, United Kingdom
B. Russell
Massachusetts Institute of Technology, Cambridge, MA 02139, USA
D. Ruterbories
University of Rochester, Rochester, NY 14627, USA
[…]
A. Saa-Hernandez
Instituto Galego de Física de Altas Enerxías, University of Santiago de Compostela, Santiago de Compostela, 15782, Spain
R. Saakyan
University College London, London, WC1E 6BT, United Kingdom
S. Sacerdoti
Université Paris Cité, CNRS, Astroparticule et Cosmologie, Paris, France
S. K. Sahoo
Indian Institute of Technology Hyderabad, Hyderabad, 502285, India
N. Sahu
Indian Institute of Technology Hyderabad, Hyderabad, 502285, India
P. Sala
Istituto Nazionale di Fisica Nucleare Sezione di Milano, 20133 Milano, Italy
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
N. Samios
Brookhaven National Laboratory, Upton, NY 11973, USA
[…]
M. C. Sanchez
Florida State University, Tallahassee, FL, 32306 USA
A. Sánchez Bravo
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
P. Sanchez-Lucas
University of Granada & CAFPE, 18002 Granada, Spain
V. Sandberg
Los Alamos National Laboratory, Los Alamos, NM 87545, USA
D. A. Sanders
University of Mississippi, University, MS 38677 USA
S. Sanfilippo
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali del Sud, 95123 Catania, Italy
D. Sankey
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
D. Santoro
Istituto Nazionale di Fisica Nucleare Sezione di Milano, 20133 Milano, Italy
N. Saoulidou
University of Athens, Zografou GR 157 84, Greece
P. Sapienza
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali del Sud, 95123 Catania, Italy
C. Sarasty
University of Cincinnati, Cincinnati, OH 45221, USA
I. Sarcevic
University of Arizona, Tucson, AZ 85721, USA
I. Sarra
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Frascati, Frascati, Roma, Italy
G. Savage
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
V. Savinov
University of Pittsburgh, Pittsburgh, PA 15260, USA
G. Scanavini
Yale University, New Haven, CT 06520, USA
A. Scaramelli
Istituto Nazionale di Fisica Nucleare Sezione di Pavia, I-27100 Pavia, Italy
A. Scarff
University of Sheffield, Sheffield S3 7RH, United Kingdom
T. Schefke
Louisiana State University, Baton Rouge, LA 70803, USA
H. Schellman
Oregon State University, Corvallis, OR 97331, USA
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
S. Schifano
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
University of Ferrara, Ferrara, Italy
P. Schlabach
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
D. Schmitz
University of Chicago, Chicago, IL 60637, USA
A. W. Schneider
Massachusetts Institute of Technology, Cambridge, MA 02139, USA
K. Scholberg
Duke University, Durham, NC 27708, USA
A. Schukraft
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
B. Schuld
University of Colorado Boulder, Boulder, CO 80309, USA
A. Segade
University of Vigo, E- 36310 Vigo Spain
E. Segreto
Universidade Estadual de Campinas, Campinas - SP, 13083-970, Brazil
[…]
C. R. Senise
Universidade Federal de São Paulo, 09913-030, São Paulo, Brazil
J. Sensenig
University of Pennsylvania, Philadelphia, PA 19104, USA
M. H. Shaevitz
Columbia University, New York, NY 10027, USA
P. Shanahan
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
P. Sharma
Panjab University, Chandigarh, 160014, India
R. Kumar
Punjab Agricultural University, Ludhiana 141004, India
K. Shaw
University of Sussex, Brighton, BN1 9RH, United Kingdom
T. Shaw
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
K. Shchablo
Institut de Physique des 2 Infinis de Lyon, 69622 Villeurbanne, France
J. Shen
University of Pennsylvania, Philadelphia, PA 19104, USA
C. Shepherd-Themistocleous
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
[…]
W. Shi
Stony Brook University, SUNY, Stony Brook, NY 11794, USA
S. Shin
Jeonbuk National University, Jeonrabuk-do 54896, South Korea
S. Shivakoti
Wichita State University, Wichita, KS 67260, USA
I. Shoemaker
Virginia Tech, Blacksburg, VA 24060, USA
D. Shooltz
Michigan State University, East Lansing, MI 48824, USA
R. Shrock
Stony Brook University, SUNY, Stony Brook, NY 11794, USA
B. Siddi
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
M. Siden
Colorado State University, Fort Collins, CO 80523, USA
J. Silber
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
L. Simard
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
J. Sinclair
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
G. Sinev
South Dakota School of Mines and Technology, Rapid City, SD 57701, USA
Jaydip Singh
University of Lucknow, Uttar Pradesh 226007, India
J. Singh
University of Lucknow, Uttar Pradesh 226007, India
L. Singh
Central University of South Bihar, Gaya, 824236, India
P. Singh
Queen Mary University of London, London E1 4NS, United Kingdom
V. Singh
Central University of South Bihar, Gaya, 824236, India
S. Singh Chauhan
Panjab University, Chandigarh, 160014, India
R. Sipos
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
C. Sironneau
Université Paris Cité, CNRS, Astroparticule et Cosmologie, Paris, France
G. Sirri
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
K. Siyeon
Chung-Ang University, Seoul 06974, South Korea
K. Skarpaas
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
J. Smedley
University of Rochester, Rochester, NY 14627, USA
E. Smith
Indiana University, Bloomington, IN 47405, USA
J. Smith
Stony Brook University, SUNY, Stony Brook, NY 11794, USA
P. Smith
Indiana University, Bloomington, IN 47405, USA
J. Smolik
Czech Technical University, 115 19 Prague 1, Czech Republic
Institute of Physics, Czech Academy of Sciences, 182 00 Prague 8, Czech Republic
M. Smy
University of California Irvine, Irvine, CA 92697, USA
M. Snape
University of Warwick, Coventry CV4 7AL, United Kingdom
E.L. Snider
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
P. Snopok
Illinois Institute of Technology, Chicago, IL 60616, USA
D. Snowden-Ifft
Occidental College, Los Angeles, CA 90041
M. Soares Nunes
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
H. Sobel
University of California Irvine, Irvine, CA 92697, USA
M. Soderberg
Syracuse University, Syracuse, NY 13244, USA
[…]
C. J. Solano Salinas
Universidad Nacional Mayor de San Marcos, Lima, Peru
Universidad Nacional de Ingeniería, Lima 25, Perú
S. Söldner-Rembold
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
S.R. Soleti
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
N. Solomey
Wichita State University, Wichita, KS 67260, USA
V. Solovov
Laboratório de Instrumentação e Física Experimental de Partículas, 1649-003 Lisboa and 3004-516 Coimbra, Portugal
W. E. Sondheim
Los Alamos National Laboratory, Los Alamos, NM 87545, USA
M. Sorel
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
[…]
J. Soto-Oton
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
A. Sousa
University of Cincinnati, Cincinnati, OH 45221, USA
K. Soustruznik
Institute of Particle and Nuclear Physics of the Faculty of Mathematics and Physics of the Charles University, 180 00 Prague 8, Czech Republic
F. Spinella
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
J. Spitz
University of Michigan, Ann Arbor, MI 48109, USA
N. J. C. Spooner
University of Sheffield, Sheffield S3 7RH, United Kingdom
K. Spurgeon
Syracuse University, Syracuse, NY 13244, USA
D. Stalder
Universidad Nacional de Asunción, San Lorenzo, Paraguay
M. Stancari
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
L. Stanco
Istituto Nazionale di Fisica Nucleare Sezione di Padova, 35131 Padova, Italy
Universtà degli Studi di Padova, I-35131 Padova, Italy
J. Steenis
University of California Davis, Davis, CA 95616, USA
R. Stein
University of Bristol, Bristol BS8 1TL, United Kingdom
H. M. Steiner
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
A. F. Steklain Lisbôa
Universidade Tecnológica Federal do Paraná, Curitiba, Brazil
[…]
J. Stewart
Brookhaven National Laboratory, Upton, NY 11973, USA
B. Stillwell
University of Chicago, Chicago, IL 60637, USA
J. Stock
South Dakota School of Mines and Technology, Rapid City, SD 57701, USA
F. Stocker
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
T. Stokes
Louisiana State University, Baton Rouge, LA 70803, USA
M. Strait
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
T. Strauss
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
L. Strigari
Texas A&M University, College Station, Texas 77840
A. Stuart
Universidad de Colima, Colima, Mexico
J. G. Suarez
Universidad EIA, Envigado, Antioquia, Colombia
J. Subash
University of Birmingham, Birmingham B15 2TT, United Kingdom
A. Surdo
Istituto Nazionale di Fisica Nucleare Sezione di Lecce, 73100 - Lecce, Italy
L. Suter
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
C. M. Sutera
Istituto Nazionale di Fisica Nucleare Sezione di Catania, I-95123 Catania, Italy
Università di Catania, 2 - 95131 Catania, Italy
K. Sutton
California Institute of Technology, Pasadena, CA 91125, USA
Y. Suvorov
Istituto Nazionale di Fisica Nucleare Sezione di Napoli, I-80126 Napoli, Italy
Università degli Studi di Napoli Federico II , 80138 Napoli NA, Italy
R. Svoboda
University of California Davis, Davis, CA 95616, USA
S. K. Swain
National Institute of Science Education and Research (NISER), Odisha 752050, India
B. Szczerbinska
Texas A&M University - Corpus Christi, Corpus Christi, TX 78412, USA
A. M. Szelc
University of Edinburgh, Edinburgh EH8 9YL, United Kingdom
A. Sztuc
University College London, London, WC1E 6BT, United Kingdom
A. Taffara
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Pisa, Pisa PI, Italy
N. Talukdar
University of South Carolina, Columbia, SC 29208, USA
J. Tamara
Universidad Antonio Nariño, Bogotá, Colombia
H. A. Tanaka
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
S. Tang
Brookhaven National Laboratory, Upton, NY 11973, USA
N. Taniuchi
University of Cambridge, Cambridge CB3 0HE, United Kingdom
A. M. Tapia Casanova
University of Medellín, Medellín, 050026 Colombia
B. Tapia Oregui
University of Texas at Austin, Austin, TX 78712, USA
A. Tapper
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
S. Tariq
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
E. Tarpara
Brookhaven National Laboratory, Upton, NY 11973, USA
E. Tatar
Idaho State University, Pocatello, ID 83209, USA
R. Tayloe
Indiana University, Bloomington, IN 47405, USA
D. Tedeschi
University of South Carolina, Columbia, SC 29208, USA
A. M. Teklu
Stony Brook University, SUNY, Stony Brook, NY 11794, USA
J. Tena Vidal
Tel Aviv University, Tel Aviv-Yafo, Israel
P. Tennessen
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
Antalya Bilim University, 07190 Döşemealtı/Antalya, Turkey
M. Tenti
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
K. Terao
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
F. Terranova
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
Università di Milano Bicocca , 20126 Milano, Italy
G. Testera
Istituto Nazionale di Fisica Nucleare Sezione di Genova, 16146 Genova GE, Italy
T. Thakore
University of Cincinnati, Cincinnati, OH 45221, USA
A. Thea
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
A. Thiebault
Université Paris-Saclay, CNRS/IN2P3, IJCLab, 91405 Orsay, France
S. Thomas
Syracuse University, Syracuse, NY 13244, USA
A. Thompson
Texas A&M University, College Station, Texas 77840
C. Thorn
Brookhaven National Laboratory, Upton, NY 11973, USA
S. C. Timm
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
E. Tiras
Erciyes University, Kayseri, Turkey
University of Iowa, Iowa City, IA 52242, USA
V. Tishchenko
Brookhaven National Laboratory, Upton, NY 11973, USA
N. Todorović
University of Novi Sad, 21102 Novi Sad, Serbia
L. Tomassetti
Istituto Nazionale di Fisica Nucleare Sezione di Ferrara, I-44122 Ferrara, Italy
University of Ferrara, Ferrara, Italy
A. Tonazzo
Université Paris Cité, CNRS, Astroparticule et Cosmologie, Paris, France
D. Torbunov
Brookhaven National Laboratory, Upton, NY 11973, USA
M. Torti
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
M. Tortola
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
F. Tortorici
Istituto Nazionale di Fisica Nucleare Sezione di Catania, I-95123 Catania, Italy
Università di Catania, 2 - 95131 Catania, Italy
N. Tosi
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
D. Totani
University of California Santa Barbara, Santa Barbara, California 93106 USA
M. Toups
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
C. Touramanis
University of Liverpool, L69 7ZE, Liverpool, United Kingdom
D. Tran
University of Houston, Houston, TX 77204, USA
R. Travaglini
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
J. Trevor
California Institute of Technology, Pasadena, CA 91125, USA
E. Triller
Michigan State University, East Lansing, MI 48824, USA
S. Trilov
University of Bristol, Bristol BS8 1TL, United Kingdom
J. Truchon
University of Wisconsin Madison, Madison, WI 53706, USA
D. Truncali
Sapienza University of Rome, 00185 Roma RM, Italy
Istituto Nazionale di Fisica Nucleare Sezione di Roma, 00185 Roma RM, Italy
W. H. Trzaska
Jyväskylä University, FI-40014 Jyväskylä, Finland
Y. Tsai
University of California Irvine, Irvine, CA 92697, USA
Y.-T. Tsai
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
Z. Tsamalaidze
Georgian Technical University, Tbilisi, Georgia
K. V. Tsang
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
N. Tsverava
Georgian Technical University, Tbilisi, Georgia
S. Z. Tu
Jackson State University, Jackson, MS 39217, USA
S. Tufanli
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
C. Tunnell
Rice University, Houston, TX 77005
J. Turner
Durham University, Durham DH1 3LE, United Kingdom
M. Tuzi
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
J. Tyler
Kansas State University, Manhattan, KS 66506, USA
E. Tyley
University of Sheffield, Sheffield S3 7RH, United Kingdom
M. Tzanov
Louisiana State University, Baton Rouge, LA 70803, USA
M. A. Uchida
University of Cambridge, Cambridge CB3 0HE, United Kingdom
J. Ureña González
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
J. Urheim
Indiana University, Bloomington, IN 47405, USA
T. Usher
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
H. Utaegbulam
University of Rochester, Rochester, NY 14627, USA
S. Uzunyan
Northern Illinois University, DeKalb, IL 60115, USA
M. R. Vagins
Kavli Institute for the Physics and Mathematics of the Universe, Kashiwa, Chiba 277-8583, Japan
University of California Irvine, Irvine, CA 92697, USA
P. Vahle
William and Mary, Williamsburg, VA 23187, USA
S. Valder
University of Sussex, Brighton, BN1 9RH, United Kingdom
G. A. Valdiviesso
Universidade Federal de Alfenas, Poços de Caldas - MG, 37715-400, Brazil
E. Valencia
Universidad de Guanajuato, Guanajuato, C.P. 37000, Mexico
R. Valentim
Universidade Federal de São Paulo, 09913-030, São Paulo, Brazil
Z. Vallari
California Institute of Technology, Pasadena, CA 91125, USA
E. Vallazza
Istituto Nazionale di Fisica Nucleare Sezione di Milano Bicocca, 3 - I-20126 Milano, Italy
J. W. F. Valle
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
R. Van Berg
University of Pennsylvania, Philadelphia, PA 19104, USA
R. G. Van de Water
Los Alamos National Laboratory, Los Alamos, NM 87545, USA
D. V. Forero
University of Medellín, Medellín, 050026 Colombia
A. Vannozzi
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali di Frascati, Frascati, Roma, Italy
M. Van Nuland-Troost
Nikhef National Institute of Subatomic Physics, 1098 XG Amsterdam, Netherlands
F. Varanini
Istituto Nazionale di Fisica Nucleare Sezione di Padova, 35131 Padova, Italy
D. Vargas Oliva
University of Toronto, Toronto, Ontario M5S 1A1, Canada
[…]
N. Vaughan
Oregon State University, Corvallis, OR 97331, USA
K. Vaziri
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Vázquez-Ramos
University of Granada & CAFPE, 18002 Granada, Spain
J. Vega
Comisión Nacional de Investigación y Desarrollo Aeroespacial, Lima, Peru
S. Ventura
Istituto Nazionale di Fisica Nucleare Sezione di Padova, 35131 Padova, Italy
A. Verdugo
CIEMAT, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas, E-28040 Madrid, Spain
S. Vergani
University College London, London, WC1E 6BT, United Kingdom
M. Verzocchi
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
K. Vetter
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
M. Vicenzi
Brookhaven National Laboratory, Upton, NY 11973, USA
H. Vieira de Souza
Université Paris Cité, CNRS, Astroparticule et Cosmologie, Paris, France
C. Vignoli
Laboratori Nazionali del Gran Sasso, L’Aquila AQ, Italy
C. Vilela
Laboratório de Instrumentação e Física Experimental de Partículas, 1649-003 Lisboa and 3004-516 Coimbra, Portugal
E. Villa
CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland
S. Viola
Istituto Nazionale di Fisica Nucleare Laboratori Nazionali del Sud, 95123 Catania, Italy
B. Viren
Brookhaven National Laboratory, Upton, NY 11973, USA
A. Vizcaya-Hernandez
Colorado State University, Fort Collins, CO 80523, USA
T. Vrba
Czech Technical University, 115 19 Prague 1, Czech Republic
Q. Vuong
University of Rochester, Rochester, NY 14627, USA
A. V. Waldron
Queen Mary University of London, London E1 4NS, United Kingdom
M. Wallbank
University of Cincinnati, Cincinnati, OH 45221, USA
J. Walsh
Michigan State University, East Lansing, MI 48824, USA
T. Walton
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
H. Wang
University of California Los Angeles, Los Angeles, CA 90095, USA
J. Wang
South Dakota School of Mines and Technology, Rapid City, SD 57701, USA
L. Wang
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
M.H.L.S. Wang
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
X. Wang
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
Y. Wang
University of California Los Angeles, Los Angeles, CA 90095, USA
K. Warburton
Iowa State University, Ames, Iowa 50011, USA
D. Warner
Colorado State University, Fort Collins, CO 80523, USA
L. Warsame
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
M.O. Wascko
University of Oxford, Oxford, OX1 3RH, United Kingdom
D. Waters
University College London, London, WC1E 6BT, United Kingdom
A. Watson
University of Birmingham, Birmingham B15 2TT, United Kingdom
K. Wawrowska
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
University of Sussex, Brighton, BN1 9RH, United Kingdom
A. Weber
Johannes Gutenberg-Universität Mainz, 55122 Mainz, Germany
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
C. M. Weber
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
M. Weber
University of Bern, CH-3012 Bern, Switzerland
H. Wei
Louisiana State University, Baton Rouge, LA 70803, USA
A. Weinstein
Iowa State University, Ames, Iowa 50011, USA
H. Wenzel
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
S. Westerdale
University of California Riverside, Riverside CA 92521, USA
M. Wetstein
Iowa State University, Ames, Iowa 50011, USA
K. Whalen
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
J. Whilhelmi
Yale University, New Haven, CT 06520, USA
A. White
University of Texas at Arlington, Arlington, TX 76019, USA
A. White
Yale University, New Haven, CT 06520, USA
L. H. Whitehead
University of Cambridge, Cambridge CB3 0HE, United Kingdom
D. Whittington
Syracuse University, Syracuse, NY 13244, USA
M. J. Wilking
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
A. Wilkinson
University College London, London, WC1E 6BT, United Kingdom
C. Wilkinson
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
F. Wilson
STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, United Kingdom
R. J. Wilson
Colorado State University, Fort Collins, CO 80523, USA
P. Winter
Argonne National Laboratory, Argonne, IL 60439, USA
W. Wisniewski
SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
J. Wolcott
Tufts University, Medford, MA 02155, USA
J. Wolfs
University of Rochester, Rochester, NY 14627, USA
T. Wongjirad
Tufts University, Medford, MA 02155, USA
A. Wood
University of Houston, Houston, TX 77204, USA
K. Wood
Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA
E. Worcester
Brookhaven National Laboratory, Upton, NY 11973, USA
M. Worcester
Brookhaven National Laboratory, Upton, NY 11973, USA
M. Wospakrik
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
K. Wresilo
University of Cambridge, Cambridge CB3 0HE, United Kingdom
C. Wret
University of Rochester, Rochester, NY 14627, USA
S. Wu
University of Minnesota Twin Cities, Minneapolis, MN 55455, USA
W. Wu
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
W. Wu
University of California Irvine, Irvine, CA 92697, USA
M. Wurm
Johannes Gutenberg-Universität Mainz, 55122 Mainz, Germany
J. Wyenberg
Dordt University, Sioux Center, IA 51250, USA
Y. Xiao
University of California Irvine, Irvine, CA 92697, USA
I. Xiotidis
Imperial College of Science Technology and Medicine, London SW7 2BZ, United Kingdom
B. Yaeggy
University of Cincinnati, Cincinnati, OH 45221, USA
N. Yahlali
Instituto de Física Corpuscular, CSIC and Universitat de València, 46980 Paterna, Valencia, Spain
E. Yandel
University of California Santa Barbara, Santa Barbara, California 93106 USA
K. Yang
University of Oxford, Oxford, OX1 3RH, United Kingdom
T. Yang
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
A. Yankelevich
University of California Irvine, Irvine, CA 92697, USA
[…]
K. Yonehara
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
T. Young
University of North Dakota, Grand Forks, ND 58202-8357, USA
B. Yu
Brookhaven National Laboratory, Upton, NY 11973, USA
H. Yu
Brookhaven National Laboratory, Upton, NY 11973, USA
J. Yu
University of Texas at Arlington, Arlington, TX 76019, USA
Y. Yu
Illinois Institute of Technology, Chicago, IL 60616, USA
W. Yuan
University of Edinburgh, Edinburgh EH8 9YL, United Kingdom
R. Zaki
York University, Toronto M3J 1P3, Canada
J. Zalesak
Institute of Physics, Czech Academy of Sciences, 182 00 Prague 8, Czech Republic
L. Zambelli
Laboratoire d’Annecy de Physique des Particules, Université Savoie Mont Blanc, CNRS, LAPP-IN2P3, 74000 Annecy, France
B. Zamorano
University of Granada & CAFPE, 18002 Granada, Spain
A. Zani
Istituto Nazionale di Fisica Nucleare Sezione di Milano, 20133 Milano, Italy
O. Zapata
University of Antioquia, Medellín, Colombia
L. Zazueta
Syracuse University, Syracuse, NY 13244, USA
G. P. Zeller
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
J. Zennamo
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
K. Zeug
University of Wisconsin Madison, Madison, WI 53706, USA
C. Zhang
Brookhaven National Laboratory, Upton, NY 11973, USA
S. Zhang
Indiana University, Bloomington, IN 47405, USA
M. Zhao
Brookhaven National Laboratory, Upton, NY 11973, USA
E. Zhivun
Brookhaven National Laboratory, Upton, NY 11973, USA
E. D. Zimmerman
University of Colorado Boulder, Boulder, CO 80309, USA
S. Zucchelli
Istituto Nazionale di Fisica Nucleare Sezione di Bologna, 40127 Bologna BO, Italy
Università di Bologna, 40127 Bologna, Italy
J. Zuklin
Institute of Physics, Czech Academy of Sciences, 182 00 Prague 8, Czech Republic
V. Zutshi
Northern Illinois University, DeKalb, IL 60115, USA
R. Zwaska
Fermi National Accelerator Laboratory, Batavia, IL 60510, USA
Abstract
The determination of the direction of a stellar core collapse via its neutrino emission is crucial for the identification of the progenitor for a multimessenger follow-up. A highly effective method of reconstructing supernova directions within the Deep Underground Neutrino Experiment (DUNE) is introduced. The supernova neutrino pointing resolution is studied by simulating and reconstructing electron-neutrino charged-current absorption on and elastic scattering of neutrinos on electrons. Procedures to reconstruct individual interactions, including a newly developed technique called “brems flipping”, as well as the burst direction from an ensemble of interactions are described.
Performance of the burst direction reconstruction is evaluated for supernovae happening at a distance of 10 kpc for a specific supernova burst flux model.
The pointing resolution is found to be 3.4 degrees at 68% coverage for a perfect interaction-channel classification and a fiducial mass of 40 kton, and 6.6 degrees for a 10 kton fiducial mass respectively. Assuming a 4% rate of charged-current interactions being misidentified as elastic scattering, DUNE’s burst pointing resolution is found to be 4.3 degrees (8.7 degrees) at 68% coverage.
Detecting neutrinos from core-collapse supernovae opens up the possibility to study the astrophysics of stellar collapse and the properties of neutrinos and their interactions. The first and only supernova neutrino detection on Earth so far was achieved with SN1987A, when a few tens of electron-antineutrino events111We note here the use of “event” to refer to an individual recorded neutrino interaction, as per standard particle physics usage. In this paper, “burst” will refer to the ensemble of events from a single core collapse. were registered [1, 2, 3]. While it was already possible to derive valuable insights from these data (e.g., Ref. [4, 5]), a high-statistics, high-resolution detection of various neutrino flavors from a future supernova will bring invaluable advancement in particle physics and astrophysics.
Information can be derived from the detected neutrino event rate, timing, energy spectrum, flavor composition, and angular distribution.
An important characteristic feature of a supernova neutrino burst is that it emerges promptly after the core bounce, preceding the related electromagnetic phenomena. As neutrinos interact only via the weak interaction, they can escape more easily in comparison to photons, allowing the neutrino signal to be observed well before the associated electromagnetic radiation (as can gravitational radiation). The Supernova Early Warning System (SNEWS) [6] is designed such that information retrieved from the supernova neutrino signal can be made available worldwide quickly, thereby facilitating the prompt detection of subsequent multi-messenger supernova-related phenomena [7, 8].
Clearly, it is highly desirable for information about the position of the supernova in the sky to be ascertained as fast as possible. Not only is the direction useful for the localization of the supernova to enable prompt measurements, but for the case when the core collapse fails to produce a bright explosion in electromagnetic radiation (e.g. black hole formation), a neutrino burst direction may help to locate a missing progenitor using archival data.
Such pointing information is available from the neutrinos themselves. One possible strategy for determining the direction to the supernova is via “triangulation” from the relative timing of neutrinos observed at different locations around the globe [9, 10, 11, 12]. However, the most promising way to achieve precision pointing is to exploit anisotropic neutrino interactions [9] for which information about the incoming neutrino direction is preserved in the final-state particle angular distribution. Water Cherenkov detectors such as Super-Kamiokande and Hyper-Kamiokande can make use of directional Cherenkov radiation to determine the supernova direction [13].
We note also that for a known source direction, directional reconstruction of final-state particles can improve the neutrino energy determination, as well as statistical classification of interaction channels with known anisotropy. Both of these can improve the extraction of physics and astrophysics from the burst.
The Deep Underground Neutrino Experiment (DUNE) is capable of a supernova burst detection among other physics goals [14, 15]. An overview of DUNE’s supernova detection capability is given in Ref. [14].
In the several tens of kilotonnes (kton) of liquid argon (LAr) volume of DUNE, charged-current interactions of electron neutrinos (CC) on 40Ar and neutrinos of all flavors scattering elastically on electrons (eES) will result in charge signals in DUNE’s time projection chamber (TPC) and scintillation light that is read out via photosensors.
In this work, both CC and eES interactions are discussed.222We focus here on directional information derived from the expected prompt burst of tens-of-MeV neutrinos, although we note that higher energy (GeV scale or more) neutrinos may follow a supernova [16, 17]. These may provide precision pointing due to both higher intrinsic lepton-neutrino collimation and better detector performance at high energy; however, they likely arrive on a long time scale (hours to years) after the supernova. Primarily eES interactions carry directional information, but the near-isotropic CC interactions have a higher cross section. A supernova burst alert from DUNE, including pointing information, will be a valuable input to SNEWS.
The overall ability for DUNE to point to a supernova using the events recorded from a core collapse depends on several factors. Two primary factors affect the direction resolution of individually recorded neutrino events. First, there are intrinsic angular spreads between the recoil directions of the final-state products and the neutrino direction. For eES interactions, which are the most important for the pointing ability, this energy-dependent spread is very well understood from weak interaction physics; for CC it is less well known. Second, the detector angular resolution smears the reconstructed direction with respect to the final-state electron direction. Detector-resolution smearing can in principle be improved with better reconstruction algorithms, although there will be an intrinsic physical limit for a given detector configuration. For this study, we use standard DUNE reconstruction techniques (detailed in Sec. IV), with some minor improvements, as well as a novel additional technique we call “brems flipping,” which provides event-by-event a modest improvement in head-tail directional disambiguation by looking at the location of the secondary particles’ tracks relative to the primary track.
While individual-event resolution on neutrino direction is relatively modest due to both physical angular spread and instrumental uncertainty, for the statistical ensemble of events in a burst the directional determination improves approximately by the inverse square root of the number of events detected. We evaluate here the overall supernova burst pointing resolution, which will enable a successful contribution from DUNE to a multi-messenger detection of a core-collapse supernova. The scope of this study, which makes use of standard offline reconstruction software, addresses only intrinsic pointing ability and does not consider latency for dissemination of pointing information.
Section II introduces the supernova neutrino-emission model that we use. In Sec. III, the DUNE detector and the expected signatures of the neutrino interaction channels in the LAr volume are described. Sec. IV provides a description of the simulation of supernova neutrino events in DUNE and the subsequently applied reconstruction algorithms including the relevant software tools. This is followed by Sec. V on the maximum likelihood method for the burst direction determination. In Sec. VI the overall performance of supernova burst direction reconstruction for DUNE is evaluated. Finally, Sec. VII summarizes the results of the study and provides a road map for potential future studies in this area.
II Supernova neutrino emission
When a massive star has depleted all of its nuclear fuel, the outgoing radiation pressure ceases to counteract the inward gravitational pull of the star, collapsing it into a compact object such as a neutron star or a black hole. During this process, 99% of the gravitational binding energy of the remnant is emitted in the form of neutrinos with energies of a few tens of MeV over a few tens of seconds. Supernova neutrinos are released in several stages during a core collapse. At the beginning of the collapse (over tens of milliseconds), neutrinos are primarily produced via electron capture () as the star undergoes neutronization. During the subsequent accretion phase (which lasts for tens to hundreds of milliseconds), more neutrinos of all flavors are created, counteracting the shock heating of the in-falling matter. Subsequently, pairs are emitted over the next tens of seconds, such that these neutrino pairs shed most of the gravitational binding energy, thereby cooling the remnant [18].
The study of supernova neutrinos is particularly useful in providing insights on varied topics in astrophysics and neutrino physics. As neutrinos are intimately involved in the collapse and subsequent explosion processes, measurements of the supernova neutrino signal allow for the examination of the complex interactions that occur during the core collapse. Additionally, supernova neutrino signals have the important feature that the initial luminosity is roughly equally divided among flavors. The subsequent flavor transitions provide information on the neutrino mass ordering, as well as insights into exotic flavor transition physics [19, 20, 21, 22].
For this study, the GKVM [23] core-collapse supernova neutrino emission model is used to describe the neutrino energy spectrum. There are known to be fairly large (up to an order of magnitude) uncertainties on the supernova neutrino flux prediction, and result both from intrinsic progenitor variance and from theoretical uncertainties. These variances will have an impact on the total number of neutrinos detected as well as on details of the neutrino flavor composition and spectra. However, the pointing capabilities of the DUNE detector are primarily sensitive to event statistics rather than to details of the model flux. We therefore do not survey different supernova models here. The selected model produces an intermediate number of eES events among a range of models. Expected event rates without energy smearing are calculated with SNOwGLoBES [24], which computes the recoil energy distribution as described in the next section. It is assumed that the supernova explosion occurs at a distance of 10 kpc from Earth, within the Milky Way for our selected model. The total expected number of events and the events per interaction channel are presented in Tab. 1. We do not study the effects of flavor transitions333Flavor transitions may affect the total number of neutrino events, but with uncertainty not exceeding the overall flux model uncertainties; spectral modulations will be a subdominant effect on pointing capabilities..
III Supernova neutrino detection at DUNE
DUNE is unique in the sense that it will be able to register large numbers of electron neutrinos, while all other experiments of similar size, e.g., Hyper-K [25] and JUNO [26], primarily detect electron antineutrinos via inverse beta decay on free protons. Immediately following the core collapse, electron-neutrino emission through neutronization dominates and insights on the neutrino mass ordering can be gained from observed differences in the neutrino flavor composition and spectra due to oscillation dynamics within the supernova [22].
We consider neutrino detection at DUNE’s far detector liquid argon time projection chamber (LArTPC). The full planned design consists of 40 kton of fiducial mass444We note this is the fiducial mass for long-baseline physics; active mass for supernova event detection could potentially be larger. of liquid argon placed into four modules. Both single-module (10 kton) and all-modules (40 kton) performance is considered in this study. We consider DUNE’s first-module design for this study. Each of the active volumes is bound by the cathode plane assembly (CPA) and the anode plane assembly (APA) and surrounded by the field cage. A uniform electric field is created between the CPA and APA, drifting ionization charges toward the APA. Three planes of sensing wires, each with a different orientation, are located at the APA. Charge depositions on these readout planes are used to reconstruct the location of the particle energy deposition in two dimensions, and the drift timing information allows for the reconstruction of the third spatial coordinate. In addition to the LArTPCs, the Photon Detection System (PDS) is used to tag neutrino events. In this study, information on the timing of the interaction is retrieved from the detection of photon flashes to be able to estimate the charge loss during drift. More details about the photon detectors can be found in [27].
The two neutrino-interaction channels considered in this study are neutrino-electron elastic scattering interactions and CC absorption interactions.
Neutral-current and anti-neutrino CC interactions are also expected [14]. However, as these channels are known to be subdominant and are at present not well understood, they are not included in this study. In the following, the nature of the dominant interaction channels is described and the simulated energy spectra and angular distributions in the detector are discussed. For eES events, the event rates are determined from SNOwGLoBES, using the GKVM supernova model. It also provides the recoil energy distribution as input to a LArSoft [28] neutrino-electron scattering event generator. For CC events, MARLEY [29] and its LArSoft interface are applied to generate the events with the correct energy distribution. Once again the GKVM supernova model is provided as input to MARLEY in this case. All simulations are done in a subset of the full DUNE far detector geometry simulating 1.6 kton of fiducial volume to reduce memory, disk, and computing time requirements.
The simulated workspace features six planes of APAs (two APAs tall) along the beam direction. For the full detector module, there are 150 APAs proposed in total, stacked two-tall and 25 along the beam direction. Because the spatial extent of the neutrino events under consideration is much smaller than the simulated workspace size, we expect the scaling down of the simulated geometry to have a negligible effect on the conclusions of the study.
Table 1: The eES and CC event rates (total for the interaction channel in bold) for DUNE within a fiducial volume of 40 kton, and for a core collapse at a distance of 10 kpc, using the GKVM model. No triggering efficiency or detector resolution effects are applied here, in contrast to [14]. Neutrino flavor transition effects beyond those included in the GKVM model are neglected. The neutrino event rate scales linearly with detector mass and with inverse-square dependence on supernova distance.
Channel
Expected Event Count
(all flavors)
325.8
155.5
67.3
55.2
47.8
3300.0
III.1 Neutrino-Electron Elastic Scattering
The most relevant interaction for directional information is when a neutrino elastically scatters off of an electron in the LAr. This interaction applies to all flavors but the largest cross section is for . A visualization of a Geant4 event is shown in Fig. 1.
Figure 1: Geant4 illustration of the energy deposition for examples of two event types: (a) CC: A 25 MeV electron neutrino is absorbed by an argon nucleus resulting in the excitation of the nucleus and the emission of an electron. (b) eES: An incoming electron neutrino of 12 MeV scatters elastically in the LAr. Charged particles are emitted in the forward direction of the primary electron. The CC primary electron tracks are on average longer than eES tracks due to the lower energies of the eES recoils; in contrast, CC electrons tend to retain most of the energy of the incoming neutrino.
Gamma tracks are not shown in the display, but red blips (representing electrons) from gamma interactions with the argon (primarily Compton scatters) can be seen.
The direction of the scattered electron is highly correlated to the direction of the neutrino. In particular, the scattering angle is described by [30]:
where is the electron kinetic energy, is the neutrino energy, and is the electron mass.
The distribution of is given by the differential cross section:
where is the Fermi coupling constant, and and are given according to neutrino flavor [30] in Tab. 2.
Table 2: Coupling strengths in the cross section of neutrino-electron scattering from Ref. [30]. The cross section of electron scattering is enhanced due to the possibility of both neutral- and charged-current interactions occurring; -electron elastic scattering is helicity-suppressed with respect to scattering.
Flavor
The event rate split into the different neutrino flavors can be found in Tab. 1. The energy distributions of the generated neutrinos and scattered electrons are shown in Fig. 2(a, blue). The distribution of the scattering angle, computed via the energy distribution and Eq. 1, is depicted in Fig. 2(b, blue) showing the sensitivity to the direction of the incoming neutrino.
DUNE is particularly sensitive to the charged-current absorption of on (CC):
Figure 2: Comparison of the cross-section-weighted energy and angular distributions of electrons for the CC events in red and the eES events in blue. The angular plots are shown as a function of the cosine of the angle between the neutrino direction and the final-state electron direction. For both types of interactions, the supernova energy spectrum is assumed for the incoming neutrinos. Top left: distribution of the energy of incoming neutrinos (continuous line) and outgoing electrons (dashed line). Top right: distribution of the cosine of the scattering angle. Bottom: two-dimensional distributions showing the relative number of events as a function of both the neutrino energy and the cosine of the scattering angle, for the two types of interactions.
The primary observable of this interaction is the emitted ; additional observables are the de-excitation products of the excited potassium nucleus in the final state. The output energy and angular distributions of this interaction are also shown in Fig. 2 in red and Fig. 1 (a) depicts an example from Geant4.
Events from CC interactions are considerably more abundant than the aforementioned eES events (about 3000 events at a core collapse distance of 10 kpc for DUNE compared to 300 for eES in Tab. 1). Yet, the direction of the electron trajectory correlates very weakly with the neutrino direction, due to the two competing nuclear transition types, Fermi and Gamow-Teller. The electrons produced by CC-induced interactions governed by Fermi transitions have an angular distribution described by , while Gamow-Teller transitions correspond to [31] with respect to the neutrino direction. For our assumed supernova neutrino energy spectrum, the angle dependence of the two contributions happen to cancel each other out almost exactly, resulting in a nearly isotropic angular distribution for CC events, as shown in the red distribution in Fig. 2(b). Figure 3 depicts the contribution of the two matrix elements for the GKVM supernova spectrum.
The pointing resolution for a supernova burst therefore depends on a precise classifier between eES and CC events as discussed in Sec. V.
We note that the Fermi/Gamow-Teller cancellation is not perfect for other assumed flux spectra (e.g. Ref. [31]), resulting in a weak anisotropy of the CC-absorption-induced electrons for many cases. Furthermore, forbidden transitions, not currently included in MARLEY, have a backward angular distribution [32] which may have an effect on the overall CC anisotropy. Currently, there are large (and not fully understood) uncertainties on the relative components of CC nuclear transition type and hence on the CC angular distribution (as well as on the total cross section [33].)
We note that while the expected CC anisotropy is relatively weak, it should be possible to extract directional information from the CC events, especially if the competing nuclear transition types can be tagged and examined separately. In principle, the nuclear transition types can be distinguished statistically using their different nuclear de-excitation product distributions. However, given the large uncertainties on the CC angular distribution555Note that these uncertainties may be addressed directly with independent laboratory measurements. and that our assumed spectrum produces the most experimentally challenging situation— a directional signal on top of a uniform CC distribution— we evaluate the pointing ability under this simple assumption and leave detailed study and use of CC anisotropies for the future.
Figure 3: The contribution of Fermi and Gamow-Teller transitions to the angular distribution for CC events, generated from MARLEY for the GVKM flux model. For this model’s neutrino spectrum, the angular correlation cancels out to a good approximation.
IV Simulation and reconstruction of supernova events
IV.1 Simulation of events in DUNE
Full-detector Monte Carlo simulations are used to determine the statistical distribution of the particles resulting from the neutrino interactions and ultimately the supernova pointing resolution.
The eES and CC event generators in LArSoft produce particles uniformly in the workspace geometry with random neutrino directions sampled from an isotropic angular distribution. Standard radioactive and detector noise models were used during the simulations [15]. The considered radioactive background sources include contaminants of the liquid argon such as 39Ar and 85Kr, radioactive contaminants from the TPC, and neutrons from the cavern walls. We expect radiological and cosmogenic background to have a relatively small effect on the pointing quality. The dominant 39Ar background of 1 Bq/kg results in only of the order of one background-induced blip expected in the spatial vicinity of a supernova event over one drift time [34]. Therefore we expect our evaluation to be robust against evolutions of the DUNE background model. Furthermore, the background will be known precisely; it can be fully characterized from the data taken near in time to the supernova burst, which will enable optimization of background mitigation for reconstruction algorithms. The simulated electronics signal process accounts for the charge deposition and drift physics, the field response of the sensing wires, electronics and digitization response of the front-end electronics, as well as the inherent electronics noise according to Ref. [34]. The Projection Matching Algorithm [35] is used for 3D track reconstruction.
An example eES event display is shown in Fig. 4. Of note, the longest track marked by ”0” (several tens of cm) with high charge deposition corresponds to the primary electron and there are as well as several shorter tracks. These tracks result from lower energy final-state particles that were produced from the interaction (such as the deexcitation products of CC events), or secondary particles created as the primary electron travels in the detector. Notably, some secondary particles have directions that correlate to the primary electron, making them useful in determining the starting direction of the primary electron track (discussed in Sec. IV.2). Additionally, charge may also be registered as a result of background radioactive decay events, such as the decay of and electronics noise. These signals are very short compared to the primary electron, are far apart from each other, and are of low amplitude, which are all characteristics that can be used for extracting signals from background.
Figure 4: Example standard DUNE event display of an eES event. The colors show a heat map of the charge deposited on each wire segment per time tick. From top to bottom, the three plots show the view of the three wire planes, U, V, and X respectively, for one APA of the LArTPC. The trajectories with numerical labels are reconstructed tracks, where “0” corresponds to the primary electron track, while ”1” and ”2” refer to the brems particles.
IV.2 Primary track identification and energy reconstruction
The reconstruction algorithm consists of several steps.
First, the energy depositions associated with the supernova neutrinos are identified. The first step in reconstructing the event is to determine which reconstructed track corresponds to the primary electron (marked with “0” in Fig. 4). The DUNE version of the projection matching algorithm as described in [36] is applied for the track reconstruction. We show in the following that the overall performance of the low-energy reconstruction is sufficient for the task at hand. Specific mitigations for the intrinsic head-tail ambiguity of the track will be described. In a noise- and background-free scenario, the track of the primary electron typically has the greatest length (see Fig. 1), as the electron deposits more charge compared to its secondary particles. However, the presence of radioactive decay particles can complicate this, as a highly energetic radioactive decay product may deposit more charge than the primary electron and create a longer track. These scenarios are rejected by examining the spatial distance between tracks. Particles associated with the neutrino interactions are clustered together, while the radioactive particles are spaced out over the total simulated LAr volume. Therefore, the primary track is selected by examining the relative position of charge depositions. Reconstructed 3D space points are sorted by their corresponding hit charges, and for the ten space points with the highest charge depositions, the distances between these space points are calculated. The two space points with the closest spatial distance between each other among those ten space points are picked and one of them is selected as the estimated reconstructed vertex position. While this location may not exactly be the true vertex, it allows the rough localization of the position inside the detector. The track with the most associated charge in the proximity of this reconstructed vertex (within a sphere of radius of 120 cm) is finally selected to be the primary electron track.
Next, the energy of the primary electron is evaluated, as the pointing resolution of an event depends on the energy of its electron. Energy reconstruction is done by summing the total charge deposited near the identified interaction vertex (with a distance cut applied at 70 cm, five times the 14 cm radiation length of liquid Ar), and mapping charge to energy using a linear relationship. This distance cut is applied to reject the energy depositions of radioactive particles far away from the interaction vertex. Charge loss due to drift is also corrected by examining the time difference between the interaction time provided by the optical detector and the TPC signal collection time. An electron lifetime of 3000 s in LAr is assumed. Comparing the reconstructed primary electron energy after drift-time correction to the true energy of the simulated particles shows a linear relationship.
IV.3 Direction reconstruction and head-tail disambiguation
Most relevant for the supernova pointing is the primary electron’s direction. The goal is to attain a accurate “reconstruction resolution”, which goes along with minimizing the sky area that must be surveyed to confidently encompass the supernova. We expect a head-tail ambiguity in LArTPC track directions as the charge drift speed towards the anode is slow compared to the propagation of particles through the TPC material. Consequently, the track’s head-tail direction cannot be determined through timing. In other words, it is not known through timing which side of the track corresponds to the origin of the track, causing the reconstructed direction to potentially be opposite of the true charged-particle track direction. This results in a bimodal distribution of angle between the true and reconstructed directon around the true and reversed directions of the primary track. The reconstruction algorithm aims to enhance the peak pointing towards the true direction.
To evaluate the performance of the reconstruction resolution, we consider detection probabilities from scanning the sky regions with the highest likelihood values first. The cosine angles between the reconstructed direction and the truth direction are collected into a histogram. The histogram is integrated from both the forward and backward directions inwards, up until the histogram drops below a certain value. This value is lowered until the integral equals 68% of the total integral of the distribution. The resolution metric “sky fraction” is defined to be the fraction of the sky that is covered by the integration. The sky fraction can be used to define an equivalent angular resolution by multiplying with 4. If the two sides of the distribution are close to fully symmetric, the sky fraction would be twice as large as the fully unambigiuous case, corresponding to an equal probability of success for finding the supernova on each side of the sky.
The head-tail ambiguity can be resolved statistically by looking at the adjacent tracks created by secondary particles. The bremsstrahlung gamma rays from the primary electron subsequently release electrons in argon atoms via Compton scattering. These secondary electrons correlate with the forward direction of the primary electron. The directional correlation is closer when the primary electron is of higher energy. The particular procedure of disambiguating the direction of the primary electron is carried out in a process called “brems flipping”.
To apply the method, a starting-point and an end-point of the primary electron track are assumed arbitrarily. The vectors from the starting point as well as from the end point of the primary track to each secondary track are determined. Subsequently, the cosine value of the angle between each vector pointing to the secondary tracks and the vector along the primary track is evaluated. The average of these cosine values is calculated.
In the next step, the assignment of the starting-point and the end-point are switched and the same calculations are carried out. Of the two vertices, the one with the larger average cosine value is finally selected as the actual starting point of the primary track. Most of the secondary particles are emitted towards the end of the primary track in a forward direction, leading to a preference for larger cosine values (corresponding to smaller angles) as can be understood from Fig. 5. The brems-flipping method is also described in Algorithm 1 (see Appendix).
Figure 5: Illustration of brems flipping: The angles between the brems particles (blue) and the primary track marked in black correspond to the actual direction of the primary electron, while the ones marked in red belong to the incorrect opposite direction. The average of the cosine of these angles is larger in the case of the correct set of angles.
IV.4 Performance of the reconstruction algorithm on single events
The performance of the reconstruction algorithm before the application of brems flipping is shown in the top plot of Fig. 6 in blue. The distribution is centered around both the parallel and anti-parallel directions with respect to the true direction, with a minor preference for the parallel direction. The reason for this preference is that the projection matching algorithm creates many track candidates at first and then merges candidates that are close to each other and have a small angle between them. If a secondary track is sufficiently close to the primary track, the merging process may combine it with the primary track, flipping the primary track to the correct orientation.
Brems flipping extends the same principle to the separately detected brems particles, meaning that their location with respect to the track indicates the direction of the primary track.
By applying brems flipping, the magnitude of the distribution in red in the anti-parallel direction is significantly decreased, confirming that the technique is a valuable tool in resolving the ambiguity in track directions.
The bottom plot of Fig. 6 shows how brems flipping reduces the sky fraction for mono-energetic electron energies. The electron energy spectrum from eES events is given as a reference for the energy regime of interest for supernova pointing. Brems flipping has a higher impact at higher primary electron energies, as those result in more secondary tracks. The successful application of brems flipping on the full supernova neutrino MC is presented in Sec. V.
Our studies show that the performance of the tracking algorithm as a function of the initial track direction with respect to the detector coordinate system is not perfectly isotropic, as expected. The particle track’s inclination with respect to the readout wire planes affects the charge deposition on the wires [35] and therefore influences the uncertainty of the reconstructed direction. The worst performance is observed along the drift direction (detector coordinate ).
Figure 7(a) demonstrates this anisotropy. For this figure, it is assumed that there are no directional ambiguities, to reveal the full impact of this anisotropy.
Figure 7(b) is created assuming the actual performance of the reconstruction algorithm including brems flipping for the head-tail disambiguation.
While the reconstruction performance varies with track direction, the figures show that the track direction ambiguity has a stronger detector anisotropy. We therefore first evaluate pointing capability assuming a uniform performance in directional reconstruction. Subsequently, we investigate the performance of the method as a function of supernova direction in detector coordinates.
Figure 6: Effectiveness of the brems-flipping algorithm: The top plot shows the bimodal distribution of the angular difference between true and reconstructed electron directions using the supernova eES electron energy spectrum, centering around the parallel () and anti-parallel () directions. Brems flipping significantly decreases the magnitude of the anti-parallel peak. The bottom plot shows the relationship between the covered sky fraction and mono-energetic electron energy. The black curve corresponds to a perfect directional disambiguation always resulting the true direction. Brems flipping performs better at higher energies, as there are more secondary tracks to reference. The electron energy spectrum from eES in gray illustrates the energies relevant to supernova pointing.
Figure 7: Pointing resolution, defined in terms of “sky fraction” as described in the text, of individual electron events as a function of the electron’s true direction in the detector coordinate system to study anisotropies in the performance. The coordinate system is defined with being the drift directions, and being approximately the beam direction. (shown vertically) and (shown horizontally) are spherical coordinates in a coordinate system for which correspond to the direction and corresponds to the direction. The supernova eES electron energy distribution is used.
(a) shows the pointing resolution given perfect track head-tail disambiguation (b) depicts the pointing resolution for the actual performance of the reconstruction algorithm including the brems-flipping algorithm.
V Maximum likelihood method for burst pointing
A maximum likelihood method [13] is used to reconstruct the direction of a supernova burst ensemble of events from the reconstructed information of individual neutrino events. For an event with a known interaction type, the probability density function (PDF) has the functional form of , where the index indicates the neutrino channel: eES or CC. and are the reconstructed energy and direction of the specific event, and is the direction of the supernova. The PDF is formulated to be a function of the reconstructed energy as well as the inner product of the supernova direction and reconstructed electron direction, , as we make the aforementioned assumption that the event direction reconstruction quality is uniform as a function of the true neutrino direction666In principle residual detector response anisotropy can be taken into account in the likelihood PDFs.. The PDFs are normalized by energy bin, as the electron energy distribution does not affect the direction likelihood. LArSoft simulations of one million events in each channel are used to determine the PDFs. The Monte Carlo samples are divided into energy bins, spanning from 0 to 40 MeV with a width of 2 MeV per bin, from 40 MeV to 70 MeV at 10 MeV per bin, and a final bin spanning from 70 MeV from 100 MeV. Varying energy bin widths are used to ensure sufficient counting statistics at high energies, where the expected numbers of events are low.
The generated PDFs are shown in Fig. 8 for both eES and CC events. For the CC channel, the low correlation between electron and neutrino directions renders the PDF mostly flat. The eES interactions demonstrate a high correlation between the supernova direction and the reconstructed direction. A bimodal distribution is seen due to the tracks’ head-tail ambiguity, but the peak for the true direction (centered around ) has a much higher amplitude than the peak around the flipped direction () thanks to brems flipping. In particular, at higher electron energies, the pointing of eES events improves both in variance around the supernova direction, as well as in track directional disambiguation (the same effect as can be observed in Figure 6). The improved performance at high energy is further demonstrated in projections of this PDF for various energy ranges shown in Figure 9.
Figure 8: PDFs of eES and CC events. The eES PDF demonstrates a high correlation between the supernova and primary electrons’ direction increasing towards higher electron energies. The CC PDF is mostly flat, as the correlation between primary electron and neutrino directions is weak.Figure 9: Several bins for the PDF of the eES interaction in Figure 8, re-binned to 10 MeV / bin. The relative magnitude of the peak near (the parallel direction) increases as energy increases because brems flipping improves at high energies. The width of the true direction peaks also decreases, indicating a decrease in directional variance.
The supernova direction can be reconstructed from the PDFs with the maximum likelihood method. Given a set of reconstructed events, the log-likelihood as a function of supernova direction is expressed as:
Here, a perfect classification between eES and CC events is assumed, . The effect of imperfect classifications is discussed later. By parameterizing using the azimuth and zenith angles, the supernova direction is reconstructed via a two-variable minimization. For the minimization, events that have reconstructed electron energies of less than 5 MeV are excluded, as they are found to reduce pointing performance due to the relatively large uncertainties in their energy and direction reconstruction. The minimization is done by a grid search. The grid search provides the likelihood function value for all directions in the sky, which is illustrated in a directional map.
The reconstructed neutrino directions associated with a typical supernova burst are shown in Fig. 10, and the output of the reconstruction procedure is shown in Fig. 11. Two local minima are 180 degrees away from each other due to the remaining ambiguity in track direction.
Figure 10: An example directional map filled with the reconstructed electron directions for a simulated supernova burst, eES events carrying the directional information are marked in blue. Event statistics are shown for a core collapse at a distance of 10 kpc, and 40 kton of fiducial mass in the detector.Figure 11: An example directional map filled with the negative log-likelihood values and confidence contours. This map is computed from the same burst shown as in Fig. 10. Reconstruction is done assuming the classification parameters of and and a successful direction reconstruction is achieved with the actual supernova direction marked with a star in the figure.
VI Performance of reconstruction algorithm on supernova simulations
Supernova bursts are simulated to evaluate the performance of the reconstruction algorithm. The simulation is carried out by randomly picking eES and CC events from a large pool of events with uniformly distributed neutrino directions. The number of events selected per burst is based on the expected number of interactions for a 10 kpc supernova, as calculated by SNOwGLoBES (see Tab. 1). For each selected event, the reconstructed direction is rotated such that all events correspond to the same neutrino direction. To account for the effect of the anisotropic direction reconstruction performance, events are only selected if their neutrino direction lies within a cone of a 10 degree opening angle of the randomly selected supernova direction. In total, 10,000 supernova bursts are simulated and reconstructed using only eES events, and a pointing resolution (as defined in the previous section) of 3.4 degrees for a fiducial mass of 40 kton (four modules, full setup) and 6.6 degrees for one module of 10 kton is achieved.
The distribution of the truth-to-reconstruction angular difference for the burst is shown in Fig. 12 for two different assumptions on event classification quality. It is worth noting that only 0.1% of the simulated bursts’ reconstructed directions differed more than 90 degrees from the supernova position thanks to the brems flipping for both classification cases. The reconstructed directions of poorly reconstructed bursts are very close to being anti-parallel to the supernova direction.
Figure 12: Distribution of the angular difference between the reconstructed and true supernova direction, for 10,000 simulated supernova bursts. The distribution is shown for both perfect event classification and for an assumed 4% misclassification of CC events as eES as described in [37]. The ranges to the right of the respective colored dashed lines correspond to the 68% confidence intervals. The few bursts with a flipped reconstructed direction are excluded from the figure. The top figure corresponds to 10 kton fiducial mass, while it is 40 kton for the bottom figure.
In reality, the capability to distinguish between eES and CC events, with the latter carrying poor pointing information, needs to be considered. The quality of a realistic event classifier can be quantified in the form of a confusion matrix:
The elements of this matrix are expressed in the form of , which describes the portion of events of interaction type that are classified as interaction type . The matrix is normalized to be independent of the expected number of events for each interaction. Assuming no loss of events due to detector inefficiency, the rows of the matrix will sum to one. The worst-case scenario would correspond to a confusion matrix with all of its elements being , which effectively represents a completely random grouping of all events into two groups.
In adopting a more realistic classifier taking into account the misidentification of events, the aforementioned reconstruction procedure is effectively unchanged, except that the PDF used in (4) no longer corresponds to , but is rather a sum of and , weighted by the expected number of events from each interaction in the classification channel, which should be known ahead of time. As pointing information is primarily accessible from eES events, only the eES classification channel is used for direction reconstruction. Subsequently, the only elements in the confusion matrix that affect the pointing resolution are and . Figure 13 shows the pointing resolution as a function of these two matrix elements. For the worst-case scenario (where both elements are equal to ), the pointing resolution is degrees. While highly precise classifiers would provide highly accurate pointing results, even a weak classifier can yield a useful pointing resolution. At , only 20% better than random selection, the pointing resolution reaches around 40 degrees, significant enough to determine the quadrant of the sky that the supernova belongs to. An optimistic estimate of a boosted-decision-tree-based classifier is established in [37] at and . With this classifier, the pointing resolution amounts to degrees (8.7 degrees) for a fiducial mass of 40 kton (10 kton). The distribution of the reconstruction angle is compared to the perfect classification case in Fig. 12. While this study provides a rough estimation of the classification capabilities, further validation work is required to understand the classification performance in the presence of noise. Furthermore, because classification performance likely depends on the energy of the event, additional improvements in pointing resolution could be made by leveraging energy regions where the classifier performs best. Moreover, in a future fast online pipeline, real-time channel tagging before the comparison of the reconstructed events to the PDF will likely improve the overall performance as well.
Figure 13: Burst pointing resolution as a function of eES true positives () and CC false negatives (). Results assuming the fiducial mass of a single far detector module (10 kton) and all four planned modules (40 kton) are shown. For each pair of values, 1000 supernova bursts are simulated to determine the pointing resolution. Contour lines for various pointing resolution angles are also shown.
The pointing resolution is shown as a function of the number of expected events in Fig. 14, along with the corresponding progenitor distance, calculated assuming the GKVM model. The pointing resolution is roughly proportional to the supernova distance, and inversely proportional to the square root of the number of expected events, as expected.
Figure 14: Burst pointing resolution as a function of the number of detected eES events (), as well as the corresponding supernova distance. The event rates are calculated assuming the GKVM model at a given distance and are for a fiducial volume of 40 kton.
The supernova pointing resolution as a function of the supernova position in the sky is also examined. Fig. 15 shows the pointing resolution as a function of the detector coordinates. As for the results shown in Fig. 7, better pointing resolution can be seen closer to the directions (with being approximately the beam direction). Note that due to the way supernova bursts are simulated (as described at the beginning of this section), the dependence of the pointing resolution estimate is smeared around the incoming neutrino direction by the event selection radius of 10 degrees – but qualitatively, the variation in pointing resolution is of the order of a few degrees.
Fig. 16 shows the pointing resolution in the equatorial celestial coordinate system (RA/Dec), as a function of declination, averaged over right ascension. The figure also depicts the expected declination distribution for galactic supernovae to illustrate the most likely directions for a supernova to occur. Notably, because the direction of the detector is positioned at around of declination, the resolution is best around the same angle.
Figure 15: Burst pointing resolution as a function of the direction of the supernova, given in the detector coordinate system ( is the drift direction, is approximately the beam direction). (shown vertically) and (shown horizontally) are spherical coordinates in a coordinate system for which corresponds to the direction and correspond to the direction. Pointing resolution is given for the fiducial volumes of 10 kton and 40 kton. Perfect classification between eES and CC events is assumed in these plots. Note that the local pointing resolution is smeared by the event selection radius of 10 degrees.Figure 16: Burst pointing resolution as a function of the supernova direction, shown as a function of declination. The pointing resolution is averaged over right ascension. Shown on a separate axis is the expected declination distribution for galactic supernovae [38].
VII Summary and outlook
An analysis framework for reconstructing supernova neutrino burst directions at DUNE is described. By reconstructing the eES and CC events during a supernova burst, the supernova direction can be determined from the correlation between the supernova neutrino and the outgoing primary electron created in the interaction with the LAr. The dominant interaction channels are CC and eES, with only the latter containing significant accessible information on the neutrino direction, but being an order of magnitude less frequent than the former.
Monte Carlo simulations in LArSoft were used to generate a data set of supernova bursts to evaluate the performance of the reconstruction algorithm for the planned DUNE detector setup.
From this data set, the energy and direction of the primary electron were reconstructed. To reduce head-tail ambiguities a highly effective technique called brems flipping was developed.
Finally, the pointing resolution is derived for ensembles of supernova burst events using a maximum likelihood method. With perfect event classification of the two event types considered, the pointing resolution is 3.4 degrees (6.6 degrees) with 68% coverage for supernovae at a distance of 10 kpc and an effective fiducial volume of 40 kton (10 kton). For a moderately optimistic classification performance, incorrect classification of 4% CC events as eES, the estimated pointing resolution is 4.3 degrees (8.7 degrees) with 68% coverage. The results presented here represent an average over all supernova directions in the sky.
In reality, there is a minor anisotropy in the burst reconstruction capability of the LAr TPC, resulting in a variation of the order of a few degrees in resolution.
A continued effort should be carried out to further improve the current supernova pointing capabilities of DUNE. Prompt dissemination of directional information will be critical for multi-messenger astronomy, and efforts are underway to enable low-latency pointing with DUNE. Furthermore, given a core collapse at the most frequently expected distance from Earth for a galactic supernova, detection of supernova neutrino events can be expected in multiple large-scale neutrino detectors. By combining DUNE’s data with events in other detectors, one may take advantage of greater event statistics as well as the strengths of different detector technologies, resulting in a noticeable improvement to the current pointing resolution. Furthermore, we anticipate that steady improvements to pattern recognition technologies, especially via machine-learning techniques, will enhance the performance of next-generation particle track reconstruction and head-tail disambiguation algorithms, allowing for additional gains in pointing ability. Extended investigation of the eES/CC classification at DUNE is required, as this currently provides the greatest uncertainty in the pointing resolution. Furthermore, subdominant interaction channels should be considered as well. As a final note, laboratory investigation of neutrino interactions in the few-tens-of-MeV range in argon will be essential to fully understand the detector directional response [39].
Acknowledgements
The ProtoDUNE-XX detector was constructed and operated on the CERN Neutrino Platform.
We gratefully acknowledge the support of the CERN management, and the
CERN EP, BE, TE, EN and IT Departments for NP04/ProtoDUNE-SP.
This document was prepared by the DUNE collaboration using the
resources of the Fermi National Accelerator Laboratory
(Fermilab), a U.S. Department of Energy, Office of Science,
HEP User Facility. Fermilab is managed by Fermi Research Alliance,
LLC (FRA), acting under Contract No. DE-AC02-07CH11359.
This work was supported by
CNPq,
FAPERJ,
FAPEG and
FAPESP, Brazil;
CFI,
IPP and
NSERC, Canada;
CERN;
MŠMT, Czech Republic;
ERDF,
H2020-EU and
MSCA, European Union;
CNRS/IN2P3 and
CEA, France;
INFN, Italy;
FCT, Portugal;
NRF, South Korea;
CAM,
Fundación “La Caixa”,
Junta de Andalucía-FEDER,
MICINN, and
Xunta de Galicia, Spain;
SERI and
SNSF, Switzerland;
TÜBİTAK, Turkey;
The Royal Society and
UKRI/STFC, United Kingdom;
DOE and
NSF, United States of America.
This research used resources of the
National Energy Research Scientific Computing Center (NERSC),
a U.S. Department of Energy Office of Science User Facility
operated under Contract No. DE-AC02-05CH11231.
Appendix A Appendix
The details of the code implementation of the descriped brems-flipping algorithm are shown in Algorithm 1.
input : Two candidate vertices with positions , and momentum directions , ;